### v1

# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

Abstract
As intelligent reflecting surface (IRS) has emerged as a new and promising technology capable of configuring the wireless environment favorably, channel estimation for IRS-assisted multiple-input multiple-output (MIMO) systems has garnered extensive attention in recent years. While various algorithms have been proposed to address this challenge, there is a lack of rigorous theoretical error analysis. This paper aims to address this gap by providing theoretical guarantees in terms of stable recovery of channel matrices for noisy measurements. We begin by establishing the equivalence between IRS-assisted MIMO systems and a compact tensor train (TT)-based tensor-on-tensor (ToT) regression. Building on this equivalence, we then investigate the restricted isometry property (RIP) for complex-valued subgaussian measurements. Our analysis reveals that successful recovery hinges on the relationship between the number of user terminals (in the uplink scenario) or base stations (in the downlink scenario) and the number of time slots during which channel matrices remain invariant. Utilizing the RIP condition, we analyze the theoretical recovery error for the solution to a constrained least-squares optimization problem, including upper error bound and minimax lower bound, demonstrating that the error decreases inversely with the number of time slots and increases proportionally with the number of unknown elements in the channel matrices. In addition, we extend our error analysis to two more specialized IRS-assisted MIMO systems, incorporating low-rank channel matrices or an unknown IRS. Furthermore, we explore a multi-hop IRS scheme and analyze the corresponding recovery errors. Finally, we introduce and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–to validate our conclusions through simulations.
Index Terms:
Intelligent reflecting surface, restricted isometry property, error analysis, tensor train-based tensor-on-tensor regression.
I Introduction
In a typical wireless propagation environment, transmitted signals undergo attenuation and scattering due to absorption, reflection, unexpected interference, diffraction, and refraction phenomena. Multipath propagation is generally acknowledged as a primary limiting factor in the performance of wireless communication systems [1]. While various physical layer techniques, such as advanced modulation/demodulation and precoding/decoding schemes, alongside the utilization of the mmWave/subTerahertz band, have been devised to counteract these adverse effects at the communication endpoints, it is evident that a plateau has been reached in terms of attainable data rates and performance reliability.
Considering the wireless environment as an additional variable for optimization holds the promise of significant performance gains. This promise is realized through the innovative concept of intelligent reflecting surface (IRS) [2, 3, 4, 5, 6, 7, 8, 9], which can dynamically reshape the wireless propagation environment to exhibit more desirable characteristics. An IRS is essentially a 2D surface composed of numerous tunable units, which can be implemented using cost-effective antennas or metamaterials and controlled in real-time to manipulate communication channels without emitting signals of their own. Recently, IRS-assisted communications have garnered significant attention [10], owing to their potential to enhance efficiency, communication range, and capacity in wireless communication systems. Nevertheless, to fulfill the role of an IRS and achieve efficient and dependable wireless communication, the acquisition of precise channel state information using channel estimation techniques is imperative, posing a formidable challenge. One particular hurdle stems from the assumption that an IRS typically comprises passive elements, thereby necessitating the receiver to estimate the cascaded channel based on pilots transmitted by the transmitter through the IRS. Recently, numerous matrix-based methods for channel estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] have been proposed and demonstrated commendable performance in practical implementation.
Building on these advancements, another promising approach leverages tensor models for channel estimation. Various tensor-based wireless communication systems, including mmWave [25, 26, 27, 28] and IRS-assisted [29, 30, 31, 32, 33] multiple-input multiple-output (MIMO) models, have been developed, providing a concise framework for channel estimation. These models aim to accommodate diverse factors like spatial, temporal, spectral, coding, and polarization diversities. Compared to matrix-based models, tensor-based models offer greater flexibility in design, allowing for more effective exploitation of the inherent tensor structure, which can lead to improved recovery accuracy. The canonical polyadic (CP) decomposition [34] was initially applied in the IRS-assisted MIMO model [29, 30, 31] for its compact representation. To convert the optimization problem into a CP factorization problem, an orthogonal pilot signal matrix is assumed, and the IRS-assisted MIMO system needs to satisfy a coefficient constraint, ensuring the uniqueness of factor identification. Due to the lack of efficient and quasi-optimal decomposition methods, the alternating least squares method is commonly employed to tackle this challenge. On the other hand, since CP decomposition combined with an identity tensor can be viewed as a special case of the Tucker decomposition [35], [33] constructed a Tucker-based IRS-assisted MIMO model and applied a quasi-optimal method–Higher Order Singular Value Decomposition–to achieve efficient channel estimation.
Although various optimization algorithms for IRS-assisted MIMO systems have demonstrated excellent performance, existing analyses of stable recovery focus solely on the required number of time slots by exploiting the uniqueness of solutions in matrix-based models [36] or the uniqueness conditions of tensor decompositions [30]. Despite these advances, there is no comprehensive error analysis for channel estimation in IRS-assisted MIMO systems, leading to the following critical question:
Question: What is the optimal recovery error for the channel estimation of the IRS-assisted MIMO system?
Our contribution:
In this paper, we affirmatively address the main question through error analyses in the tensor-based muliple-access/multi-user IRS-assisted MIMO system. Instead of using CP or Tucker decompositions, we first rigorously establish that the combination of channel matrices and IRS is equivalent to the tensor train (TT) format [37], which has been extensively utilized in quantum physics for efficiently representing quantum states [38, 39, 40]. Additionally, we demonstrate that the IRS-assisted MIMO model aligns with the tensor-on-tensor (ToT) regression model [41, 42, 43] without any assumptions regarding the pilot signals. This mode extends the classical matrix/tensor regression framework, in which the response is the vector [44, 45, 46, 47, 48, 49], to accommodate transmitted signals (represented by the second word “tensor” in the name) and the received signals (represented by the second word “tensor” in the name) as tensors. Building on the definitions outlined above, the channel estimation in a standard IRS-assisted MIMO model can be equivalently viewed as a recovery problem within the TT-based ToT regression model. Importantly, while other matrix/tensor-based models can be equivalently transformed into this framework, the TT-based ToT regression model is particularly advantageous because its canonical TT format [50, Theorem 1] facilitates theoretical analysis and the ToT model provides a compact representation that streamlines the design of optimization methods tailored to various pilot signal configurations. Consequently, this paper primarily focuses on analyzing this model.
Building on the previous discussion, our problem is reformulated as a recovery problem within the TT-based ToT regression model, representing a multiple-access/multi-user MIMO communication system assisted by an IRS in an uplink communication scenario. Towards that goal, we initially establish the restricted isometry property (RIP) for complex-valued subgaussian measurements, where each pilot signal element is a complex-valued subgaussian random variable. This analysis reveals an optimal size relationship between user terminals (UTs) and time slots. Notably, for a downlink communication scenario, RIP relates to base stations (BSs) and time slots. Subsequently, we investigate channel matrices recovery of the least square optimization problem from subgaussian measurements and derive recovery bounds relative to the number of time slots. Our findings illustrate that recovery error is inversely proportionate to the number of time slots, and successful recovery depends on the relationship between UTs and time slots. Aside from aforementioned constraints, our recovery analyses remain free from other coefficient constraints. Moreover, we extend recovery error considerations to include low-rank channel matrices [30] and unknown IRS [51]. Next, we explore a multi-hop IRSs scheme and analyze corresponding recovery errors, laying the groundwork for future multi-hop IRSs communication design. Finally, we propose two nonconvex optimization algorithms–alternating least squares and alternating gradient descent and simulations affirm the validity of theoretical analyses.
Notations
We use calligraphic letters (e.g., $\mathcal{Y}$) to denote tensors, bold capital letters (e.g., $\bm{Y}$) to denote matrices, except for $\bm{X}_{i}$ which denotes the $i$-th order-$3$ tensor factors in the TT format ($i=2,\dots,N+M-1$), bold lowercase letters (e.g., $\bm{y}$) to denote vectors, and italic letters (e.g., $y$) to denote scalar quantities. $a^{*}$ is the complex conjugate of $a$. Elements of matrices and tensors are denoted in parentheses, as in Matlab notation. For example, $\mathcal{X}(s_{1},s_{2},s_{3})$ denotes the element in position $(s_{1},s_{2},s_{3})$ of the order-3 tensor $\mathcal{X}$. The inner product of $\mathcal{A}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ can be denoted as $\langle\mathcal{A},\mathcal{B}\rangle=\sum_{s_{1}=1}^{d_{1}}\cdots\sum_{s_{N}=1}^{d_{N}}\mathcal{A}(s_{1},\dots,s_{N})\mathcal{B}(s_{1},\dots,s_{N})$. $\|\mathcal{X}\|_{F}=\sqrt{\langle\mathcal{X},\mathcal{X}\rangle}$ is the Frobenius norm of $\mathcal{X}$. Tensor contraction is defined as $\mathcal{A}\times_{i}^{j}\mathcal{B}$ of size $d_{1}\times\cdots\times d_{i-1}\times\times d_{i+1}\times\cdots\times d_{N}\times h_{1}\times\cdots\times h_{j-1}\times h_{j+1}\times\cdots\times h_{M}$ with $(s_{1},\dots,s_{i-1},s_{i+1},\dots,s_{N},f_{1},\dots,f_{j-1},f_{j+1},\dots,f_{M})$-th entry being $\sum_{k}\mathcal{A}(s_{1},\dots,s_{i-1},k,s_{i+1},\dots,s_{N})\mathcal{B}(f_{1},\dots,\\ f_{j-1},k,f_{j+1},\dots,f_{M})$ for $\mathcal{A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{i-1}\times d_{k}\times d_{i+1}\times\cdots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{h_{1}\times\cdots\times h_{j-1}\times d_{k}\times h_{j+1}\times\cdots\times h_{M}}$. The procedure of $\mathcal{A}\times_{i_{1},\dots,i_{n}}^{j_{1},\dots,j_{n}}\mathcal{B}$ can be viewed as a sequence of $n$ tensor contractions $\mathcal{A}\times_{i_{k}}^{j_{k}}\mathcal{B},k\in[n]$. $\|\bm{X}\|$ represents the spectral norm of the matrix $\bm{X}$ For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b=\Omega(a)$ indicates that $b\geq ca$ for some universal constant $c$.
II Single-Hop IRS-Assisted MIMO System
Model formulation
We consider a muliple-access/multi-user MIMO communication system assisted by an IRS, in which the $P$ base stations (BSs) receive the signals transmitted by $U$ user terminals (UTs) via the IRS. The terminology employed in this paper is based on an uplink communication scenario, where the transmitter is referred to as the UT and the receiver as the BS. However, it is important to note that our signal model and theoretical analysis can be equally applied to the downlink case by merely reversing the roles of the transmitter and the receiver. Without loss of generality, we assume that each BS and each UT are equipped with the same number of antennas, denoted as $L$ and $M$ respectively. The IRS is composed of $N$ passive elements, or unit cells, capable of individually adjusting their reflection coefficients (i.e., phase shifts). Due to unfavorable propagation conditions, BS-UT channels are neglected [29]. We also assume quasi-static flat fading channel model, where all the channels remain invariant during $T$ time slots. Assuming a block-fading channel, the received signal model is usually given as follows [15, 29, 30, 31]
$\bm{y}(t)=\bm{H}\text{diag}(\bm{s}(t))\bm{G}\bm{x}(t)+\bm{w}(t),\ t\in[T],$ (1)
where the IRS-BS channel is $\bm{H}=\begin{bmatrix}\bm{H}_{1}^{\top}\cdots\bm{H}_{P}^{\top}\end{bmatrix}^{\top}\in\mathbb{C}^{LP\times N}$ in which $\bm{H}_{p}\in\mathbb{C}^{L\times N},p\in[P]$ is the $p$-th IRS-BS channel, and the UT-IRS channel is $\bm{G}=\begin{bmatrix}\bm{G}_{1}\cdots\bm{G}_{U}\end{bmatrix}\in\mathbb{C}^{N\times UM}$ in which $\bm{G}_{u}\in\mathbb{C}^{N\times M},u\in[U]$ is $u$-th the UT-IRS channel. $\bm{x}(t)=\begin{bmatrix}\bm{x}_{1}^{\top}(t)\cdots\bm{x}_{U}^{\top}(t)\end{bmatrix}^{\top}\in\mathbb{C}^{UM\times 1}$ is the vector containing the transmitted pilot signals at time $t$, where $\bm{x}_{u}(t)\in\mathbb{C}^{M\times 1}$ is the $u$-th user pilot vector. The phase shift vector at time $t$ is defined as $\bm{s}(t)=\begin{bmatrix}s_{1}(t)e^{j\phi_{1}(t)}\cdots s_{N}(t)e^{j\phi_{N}(t)}\end{bmatrix}^{\top}\in\mathbb{C}^{N\times 1}$, where $s_{n}(t)\in\{0,1\}$ and $e^{j\phi_{n}(t)}\in(0,2\pi]$ represent the on or off state and the phase shift of the $n$-th IRS element at time instant $t$, respectively. $\bm{w}(t)\sim\mathcal{C}\mathcal{N}({\bm{0}},\gamma^{2}{\bf I}_{LP})$ denotes the complex additive white Gaussian noise (AWGN).
The channel coherence time $T_{s}$ is divided into $K$ blocks, where each block has $T$ time slots so that $T_{s}=KT$ [29]. Let us define $\bm{y}(k,t)=\bm{y}((k-1)T+t)$ as the received signal at the $t$-th time slot of the $k$-th block, where $t\in[T]$ and $k\in[K]$. In addition, we denote $\bm{x}(k,t)$ and $\bm{s}(k,t)$ as the pilot signal and phase shift vectors associated with the $t$-th time slot of the $k$-th block. In each block, the user transmits the same pilot signal and the IRS phase shift vector is invariant, but the phase shift vector between blocks is changeable, which means that the IRS has $K$ phase configurations. Consequently, we have $\bm{s}(k)=\bm{s}(k,t)$ and $\bm{x}(t)=\bm{x}(k,t)$. The received signal is given by
$\bm{y}(k,t)\!=\!\bm{H}\text{diag}(\bm{s}(k))\bm{G}\bm{x}(t)+\bm{w}(k,t),\ t\in[T]\ \text{and}\ k\in[K].$
By respectively stacking $\bm{y}(k,t)$, $\bm{x}(t)$ and $\bm{w}(k,t),t\in[T]$ into matrices $\bm{Y}(k)=\begin{bmatrix}\bm{y}(k,1)\cdots\bm{y}(k,T)\end{bmatrix}\in\mathbb{C}^{LP\times T}$, $\bm{X}=\begin{bmatrix}\bm{x}(1)\cdots\bm{x}(T)\end{bmatrix}\in\mathbb{C}^{UM\times T}$ and $\bm{W}(k)=[\bm{w}(k,1)\cdots$ $\bm{w}(k,T)]\in\mathbb{C}^{LP\times T}$, we have
$\bm{Y}(k)=\bm{H}\text{diag}(\bm{s}(k))\bm{G}\bm{X}+\bm{W}(k).$ (2)
The objective is to estimate the channel matrices $\bm{H}$ and $\bm{G}$ from the received signals $\{\bm{Y}(k)\}_{k=1}^{K}$ by appropriately designing the pilot signal matrix $\bm{X}$ and the phase shift vectors $\{\bm{s}(k)\}_{k=1}^{K}$. Notably, this formulation highlights $(i)$ the flexibility in designing $\bm{X}$ to optimize estimation performance, and $(ii)$ the desire to minimize the number of time slots $T$ to enhance efficiency.
Previous studies [29, 30, 52, 13, 31] have reformulated (2) as a noisy CP factorization problem. This is achieved either by treating $\bm{G}\bm{X}$ as a single matrix [29, 30, 52] under the assumption that the pseudoinverse of $\bm{X}$ exists, or by assuming that the pilot signal matrix $\bm{X}$ is orthogonal [13, 31], such as employing the discrete Fourier transform (DFT) matrix, allowing (2) to be rewritten as $\bm{Y}(k)\bm{X}^{\rm H}=\bm{H}\text{diag}(\bm{s}(k))\bm{G}+\bm{W}(k)\bm{X}^{\rm H}$. However, these assumptions impose constraints on the design of the pilot signal matrix, limiting their applicability to general cases. Moreover, the CP factorization framework used for single-hop IRS-assisted MIMO systems does not readily extend to multi-hop configurations, presenting additional challenges.
Compact tensor representation
To accommodate a more general setting for the pilot signal matrix $\bm{X}$, we can reinterpret (2) as a tensor-on-tensor (ToT) regression model. Specifically, by aggregating the received data over $K$ blocks, we can define three tensors as $\mathcal{Y}\in\mathbb{C}^{LP\times K\times T}$, $\mathcal{S}\in\mathbb{C}^{N\times K\times N}$ and $\mathcal{W}\in\mathbb{C}^{LP\times K\times T}$, where $\mathcal{Y}(:,k,:)=\bm{Y}(k)$, $\mathcal{S}(:,k,:)=\text{diag}(\bm{s}(k))$ and $\mathcal{W}(:,k,:)=\bm{W}(k)$. Then, the ground-truth channel tensor $\mathcal{B}^{\star}\in\mathbb{C}^{LP\times K\times UM}$ can be defined as follows:
[…]
Compared to directly analyzing (2), the tensor formulation enables the modeling of correlations in the signal model (2) across different phase configurations indexed by $k$. Thus, each element of $\mathcal{Y}$ can be represented as
$\mathcal{Y}(p,k,t)=\sum_{m=1}^{UM}\mathcal{B}^{\star}(p,k,m)\bm{X}(m,t)+\mathcal{W}(p,k,t).$ (4)
In this context, $\mathcal{Y}$ and $\bm{X}$ represent the first and second tensors in the “tensor-on-tensor” regression model  [^1] Without loss of generality, we can regard the matrix $\bm{X}$ as an order-$2$ tensor., respectively. It is evident that the pilot signal (measurement operator) $\bm{X}$ exclusively affects UTs, as the received signal $\mathcal{Y}$ is a tensor of size $LP\times K\times T$ and the channel tensor $\mathcal{B}^{\star}$ in (4) is a tensor of size $LP\times K\times UM$, implying the total number of antennas in the UTs $UM$ should be contingent on the number of time slots $T$.
Notice that, unlike the construction of CP decomposition, which requires transforming (3) as demonstrated in [29, 30, 52, 13, 31], (3) is inherently represented in the standard tensor train (TT) format as each element of $\mathcal{B}^{\star}$ can be expressed in a matrix product form (a detailed introduction to the TT decomposition is provided in Appendix A). By applying the tensor contraction operation [53], the TT form of $\mathcal{B}^{\star}$ can be expressed as:
$\mathcal{B}^{\star}=[\bm{H},\mathcal{S},\bm{G}]=\bm{H}\times_{2}^{1}\mathcal{S}\times_{3}^{1}\bm{G}\in\mathbb{C}^{LP\times K\times UM},$ (5)
where $(\bm{H}\times_{2}^{1}\mathcal{S})(p,k,i)=\sum_{n=1}^{N}\bm{H}(p,n)\mathcal{S}(n,k,i)$ and $(\mathcal{C}\times_{3}^{1}\bm{G})(p,k,m)=\sum_{n=1}^{N}\mathcal{C}(p,k,n)\bm{G}(n,m)$.
Therefore, we can rewrite (4) as a TT-based ToT regression model.
[…]
where (×$(\mathcal{B}^{\star}\times_{3}^{1}\bm{X})(p,k,t)=\sum_{j=1}^{UM}\mathcal{B}^{\star}(p,k,j)\bm{X}(j,t)$. The linear map $\mathcal{X}(\mathcal{B}^{\star}):\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ models the measurement process.
We now consider minimizing the following constrained least squares objective to recover the channel tensor $\mathcal{B}^{\star}$:
^$\widehat{\mathcal{B}}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\frac{1}{T}\|\mathcal{X}(\mathcal{B})-\mathcal{Y}\|_{F}^{2},$ (7)
where we define a set of order-$3$ TT format tensor as following:
[…]
$\text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}.$
The advantage of the least-squares formulation is that it allows the application of well-established tools from standard matrix/tensor regression models, such as the Restricted Isometry Property (RIP) condition [54, 55, 56, 57, 40], as detailed in the following discussion. Subsequently, we provide a rigorous theoretical analysis, deriving error bounds for channel estimation.
Error analysis
One advantage of viewing the channel estimation as the regression model is that analogous to the analysis in standard matrix/tensor regression models, a larger number of time slots $T$ tends to enhance denoising effectiveness since the error bound of channel estimation in the IRS-assisted MIMO model should heavily rely on $T$. Recently, we analyzed the statistical guarantee of the TT-based ToT regression model in [58]. However, these results cannot be directly applied to the IRS-assisted MIMO model because recovery errors pertain to the entire tensor recovery. Since the phase shift tensor $\mathcal{S}$ is often known, only the channel matrices are affected by noise. Consequently, the recovery error should be primarily related to the channel matrices. Next, we will establish the upper bound and the minimax lower bound of the recovery error for channel estimation.
To facilitate the recovery of channels $\bm{H}$ and $\bm{G}$ from their linear measurements $\mathcal{Y}$, the pilot signal (regression/measurement operator) need adhere to specific properties. One such desirable property is the RIP, extensively explored and popularized in compressive regression literature [54, 55, 56, 59]. This property has also been extended to structured tensors [57, 40].
Theorem 1.
Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable  [^2] Note that a complex-valued random variable $X$ is subgaussian if and only if its both real part $\operatorname{Re}\left\{X\right\}$ and imaginary part $\operatorname{Im}\left\{X\right\}$ are real subgaussian random variables. Here are some classical examples of subgaussian distributions. (Gaussian) A standard complex Gaussian random variable $X=\operatorname{Re}\left\{X\right\}+i\operatorname{Im}\left\{X\right\}$ with $\operatorname{Re}\left\{X\right\}$ and $\operatorname{Im}\left\{X\right\}$ being independent and following $\mathcal{N}(0,\frac{1}{2})$, is a subgaussian random variable. (Bernoulli) A Bernoulli random variable $X$ that takes values $-1$ and $1$ with equal probability is a subgaussian random variable. (Digital modulation signals) The phase shift keying (PSK) [60], quadrature amplitude modulation (QAM) [61] and pulse amplitude modulation (PAM) [62] exhibit subgaussian random variable properties. . Let $\delta_{UM}\in(0,1)$ be a positive constant. Then, for any channel tensor $\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}$, when the number of time slots satisfies
$T\geq C\cdot\frac{UM}{\delta_{UM}^{2}},$ (9)
with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
$(1-\delta_{UM})\|\mathcal{B}\|_{F}^{2}\leq\frac{1}{T}\|\mathcal{X}(\mathcal{B})\|_{F}^{2}\leq(1+\delta_{UM})\|\mathcal{B}\|_{F}^{2},$ (10)
where $c$ and $C$ are positive constants.
The proof is presented in Appendix B. Theorem 1 guarantees the RIP for subgaussian measurement ensembles, with the number of time slots $T$ scaling linearly only in relation to the total number of antennas ($UM$) in the UTs  [^3] In a downlink communication scenario, the parameter $T$ will be connected to the total number of antennas ($LP$) in the BSs.. Subgaussian random variables, including Gaussian, Bernoulli, PSK, QAM, and PAM signals, can be conveniently designed as pilot signals in practical applications. Notably, while each element of the Haar-distributed random unitary matrix is a subgaussian random variable, the same does not hold for the elements of the normalized DFT matrix, $\frac{1}{\sqrt{T}}\bm{X}$, a commonly employed pilot signal matrix. Nevertheless, the DFT matrix satisfies (10) when $T\geq UM$, rendering the subsequent conclusions with $\delta_{UM}=0$ applicable to them as well. Our analysis centers on the subgaussian case to streamline the discussion. When RIP holds, then for any two distinct TT format channel tensors $\mathcal{B}_{1},\mathcal{B}_{2}\in\mathbb{B}_{N,\mathcal{S}}$, noting that $\mathcal{B}_{1}-\mathcal{B}_{2}\in\mathbb{B}_{2N,\mathcal{S}}$ is also a TT format channel tensor according to (A) in Appendix A, we have distinct measurements since
[…]
which guarantees the possibility of exact recovery of channel matrices in the channel tensor.
Next, we present a formal analysis of the upper bound for ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ with ^$\widehat{\mathcal{B}}\in\mathbb{B}_{N,\mathcal{S}}$. Here, we provide a concise summary of the main result.
Theorem 2.
(Upper bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}$ in (8) where $\bm{H}$ and $\bm{G}$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(LPN+UMN)}$ for a positive constant $c_{1}$, we have
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (12)
where $\widehat{\mathcal{B}}$ is the solution to (7).
The detailed analysis is provided in Appendix C. Theorem 2 guarantees a stable recovery of the ground truth $\mathcal{B}^{\star}$ when the number of time slots $T$ is linearly proportionate to $LPN+UMN$. Combining the condition $T\geq\Omega(UM/\delta_{UM}^{2})$ in (9) with ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
$T\geq\max\bigg{\{}\Omega\bigg{(}\frac{UM}{\delta_{UM}^{2}}\bigg{)},\Omega\bigg{(}\frac{(1+\delta_{UM})(LPN+UMN)\gamma^{2}}{(1-\delta_{UM})^{2}\epsilon^{2}}\bigg{)}\bigg{\}}$
to ensure a small recovery error.
Next we consider two special cases: $(i)$ Theorem 2 assumes that the channel matrices $\bm{H}$ and $\bm{G}$ are full rank, but in some scenarios, such as millimeter-wave MIMO systems, the presence of a large number of transmit/receive antennas combined with scattering-poor propagation may lead to low-rank channel matrices $\bm{H}$ and $\bm{G}$ [63, 15]. When the signal travels between the BS and IRS across $r_{1}$ clusters, and between the IRS and the UT across $r_{2}$ clusters, it is reasonable to assume that $\bm{H}$ is a rank-$r_{1}$ matrix and $\bm{G}$ is a rank-$r_{2}$ matrix. By leveraging the covering number of low-rank matrices as introduced in [64, Lemma 3.1] and applying it to (40) in Appendix C, we can readily extend (12) to
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$
$\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})((LP+N)r_{1}+(UM+N)r_{2})}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (13)
where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N}\ \text{with}\ \text{rank}(\bm{H})$
$=r_{1},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{with}\ \text{rank}(\bm{G})=r_{2},\text{are unknown and}$
$\mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}$. It is important to note that a comparison between (12) and (II) reveals that the improvement in recovery error due to low-rankness occurs only when $r_{1}\leq\frac{LPN}{LP+N}$ and $r_{2}\leq\frac{UMN}{UM+N}$.
$(ii)$ It is crucial to emphasize that the elements of the IRS in outdoor scenarios, being exposed to weather and atmospheric conditions, may encounter unknown blockages and time-dependent fluctuations in their phase and amplitude responses [51]. Consequently, the IRS tensor $\mathcal{S}$ deviates from its intended structure, and the assumption of perfect knowledge of all phase shifts at the receiver may not be valid in such cases. To analyze the scenario where $\mathcal{S}$ is unknown, by introducing the covering number for $\mathcal{S}$ to the proof in Appendix C , we can obtain
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\big{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+KN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\big{)},$
where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{are unknown}\}$. Apart from the additional term $KN$, this upper bound closely aligns with (12). Hence, we omit redundant discussions and concentrate exclusively on the scenario where $\mathcal{S}$ is perfectly known.
Our second goal is to establish a minimax lower bound using Gaussian ensemble design. This lower bound proves to be a valuable tool for gaining insights into the parameters that influence the achievable error of a given problem. Furthermore, it reveals that the upper bound in (12) is tight and optimal. Following the derivation in Appendix D, we arrive at:
Theorem 3.
(Minimax lower bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) In the context of the TT-based ToT regression model in (6), we assume a channel tensor $\mathcal{B}^{\star}$ from (8), where $\bm{H}$ and $\bm{G}$ are full rank and $N\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg{)}.$ (14)
Note that the constraint $N\geq C^{\prime}$ is not strictly necessary, primarily simplifying the theoretical analysis in Appendix D.
III Generalization to Multi-Hop IRS-Assisted MIMO System
The preceding section focuses on the single-hop IRS assisted MIMO system, recognized as a promising technology to address the propagation distance challenge [65, 6, 66, 67, 68]. However, in specific scenarios like urban areas or satellite-to-indoor communications, employing a multi-hop IRSs scheme becomes imperative to surmount severe signal blockage between BSs and UTs for enhanced service coverage [69, 70, 71]. Furthermore, in the context of Terahertz (0.1-10 THz) band communication, considered a promising technology for achieving ultra-high speed and low-latency communications, deploying multiple passive IRSs between BSs and UTs proves effective in overcoming inherent propagation attenuations.
Tensor-based model formulation
In this section, we explore the $D$-IRS scheme connecting UTs and BSs, assuming unavailability of channels from UT to BS, from UT to the $i$-th IRS ($1<i\leq D$), from the $j$-th IRS ($1\leq j<D$) to BS, and between non-adjacent IRSs due to blockage or excessive path loss. In detail, we denote the UT-$1$-th IRS channel as $\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM}$, the $D$-th IRS-BS channel as $\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}}$, the channel from the $d$-th IRS to the $(d+1)$-th IRS as $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ for $d\in[D-1]$, and represent the $d$-th IRS as $\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}}$ for $d\in[D]$. Here, $\mathcal{S}_{d}(:,k,:)=\text{diag}(\bm{s}_{d}(k))\in\mathbb{C}^{N_{d}\times N_{d}}$, and the design of $\bm{s}_{d}(k)$ is same with (2). We define each element of the channel tensor $\mathcal{B}^{\star}_{D}\in\mathbb{C}^{LP\times K\times UM}$ as follows:
[…]
$\cdot\mathcal{S}_{1}(:,k,:)\bm{B}_{0}(:,j).$
Assume that each element of $\mathcal{W}$ is subject to complex additive white Gaussian noise, the received signal is given by
$\widetilde{\mathcal{Y}}(p,k,t)\!=\!\sum_{j=1}^{UM}\mathcal{B}^{\star}_{D}(p,k,j)\bm{X}(j,t)\!+\!\mathcal{W}(p,k,t),$ (16)
where, in contrast to the multi-hop scheme [69], we assume the absence of a direct channel from the BS to the UT to streamline the analysis.
Note that each element $\mathcal{B}^{\star}_{D}(p,k,j)$ can still be conceptualized as an element of an order-$2D+1$ TT format tensor due to its multiplication form resembling (29). Specifically, to establish a connection between the channel tensor $\mathcal{B}^{\star}_{D}$ and the TT format tensor, we can reinterpret $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ as an order-$3$ tensor $\widetilde{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times 1\times N_{d}}$. Accordingly, we define ^$\widehat{\mathcal{B}}_{D}^{\star}$ as the result of a sequence of tensor contractions  [^4] Here, the tensor contraction operation $\mathcal{A}_{1}\times_{d}^{1}\mathcal{A}_{2}$ for $\mathcal{A}_{1}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ and $\mathcal{A}_{2}\in\mathbb{C}^{N_{1}\times M_{2}\times M_{3}}$ results in a new tensor $\mathcal{A}_{3}$ of size $N_{1}\times\cdots\times N_{d-1}\times N_{d+1}\times\cdots\times N_{D}\times M_{2}\times M_{3}$, with the $(n_{1},\dots,n_{d-1},n_{d+1},\dots,n_{D},m_{2},m_{3})$-th element being $\sum_{n_{d}}\mathcal{A}_{1}(n_{1},\dots,n_{d})\mathcal{A}_{2}(n_{d},m_{2},m_{3})$.: $\bm{B}_{D}\times_{2}^{1}\mathcal{S}_{D}\times_{3}^{1}\widetilde{\bm{B}}_{D-1}\times_{4}^{1}\cdots\times_{2D-1}^{1}\widetilde{\bm{B}}_{1}\times_{2D}^{1}\mathcal{S}_{1}\times_{2D+1}^{1}\bm{B}_{0}$. As a result, ^$\widehat{\mathcal{B}}_{D}^{\star}$ can be represented as an order-$2D+1$ TT format tensor with dimensions $LP\times K\times 1\times K\times\cdots\times 1\times K\times UM$ with TT ranks  [^5] The definition of TT ranks is provided in Appendix A. $(N_{D},N_{D},N_{D-1},N_{D-1},\dots,N_{1},N_{1})$. Further insights can be gained, revealing that
$\mathcal{B}^{\star}_{D}(p,k,j)$
={^$=\begin{cases}\widehat{\mathcal{B}}_{D}^{\star}(p,k_{1},1,k_{2},1,\dots,k_{D},j),&k_{1}=\dots=k_{D}=k,\\ 0,&\text{otherwise},\end{cases}$ (17)
where it becomes evident that $\mathcal{B}^{\star}_{D}$ emerges as a sampled outcome of ^$\widehat{\mathcal{B}}_{D}^{\star}$. To streamline the notation, we introduce the notation $\mathcal{B}^{\star}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}$. Consequently, we define the TT-based ToT regression model as following:
[…]
Now, we consider the following constrained least squares objective:
^$\widehat{\mathcal{B}}_{D}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})-\widetilde{\mathcal{Y}}\|_{F}^{2},$ (19)
where we define the set of $\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$ as follows:
[…]
$\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},$
$d\in[D-1]\ \text{are unknown and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},$
$d\in[D]\ \text{are known}\},$
in which $N=\max_{d}N_{d}$. Note that each element of $\mathcal{B}_{D}$ follows (15). In the subsequent section, we present a comprehensive error analysis of the channel estimation process.
Error analysis
Before presenting the error analysis, we first ensure that the measurements satisfy the RIP condition. By following a similar derivation as provided in Appendix B, the RIP condition for $\mathcal{X}(\mathcal{B}_{D})$ can be directly obtained.
Theorem 4.
Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable. Let $\delta_{UM}\in(0,1)$ be a positive constant. Then, for any channel tensor $\mathcal{B}^{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$, when the number of time slots satisfies
$T\geq C\cdot\frac{UM}{\delta_{UM}^{2}},$ (21)
with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
$(1-\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2}\!\leq\!\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})\|_{F}^{2}\!\leq\!(1+\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2},$ (22)
where $c$ and $C$ is a universal constant.
Similar to Theorem 1, Theorem 4 ensures that the number of time slots is solely contingent upon the total number of antennas (UM) in the UTs. This is attributed to the fact that the pilot signal matrix $\bm{X}$ is exclusively associated with the UT-$1$-th IRS channel $\bm{B}_{0}$. Then we can further analyze the recovery error ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$.
Theorem 5.
(Upper bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}_{D}$ in (20) in which $\bm{B}_{d},d=0,\dots,D$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}$ for a positive constant $c_{1}$, we have
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (23)
where $N_{0}=UM$, $N_{D+1}=LP$ and  ^$\widehat{\mathcal{B}}_{D}$ is the solution to (19).
The proof has been provided in Appendix E. Theorem 5 ensures an optimal recovery error bound, given that the degree of freedom for the unknown variables in (19) is $O(LPN_{D}+UMN_{1}+\sum_{i=1}^{D-1}N_{d}N_{d+1})$. However, it is noteworthy that while increasing the number of IRSs can expand service coverage, the recovery error of channel estimation may also increase when the number of time slots $T$ is limited. Therefore, in practical implementation, a balance should be struck between service coverage and recovery error. Similar to the analysis of the minimax lower bound using the Gaussian ensemble design in Theorem 3, this result can also be extended to multi-hop IRS-assisted MIMO systems.
Theorem 6.
(Minimax lower bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) In the context of the TT-based ToT regression model in (18), we assume a channel tensor $\mathcal{B}^{\star}_{D}$ from (20), where $\bm{B}_{d},d=0,\dots,D$ are full rank and $\min N_{d}\geq C^{\prime}$ with $C^{\prime}$ being a universal constant. Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
$\inf_{\widehat{\mathcal{B}}_{D}}\sup_{\mathcal{B}^{\star}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{\sum_{d=0}^{D}N_{d}N_{d+1}}{T}}\gamma\bigg{)}.$ (24)
Finally, akin to the analysis of a single-hop IRS-assisted MIMO system, it is imperative to address the following two special cases. $(i)$ First, when $\bm{B}_{d},d=0,\dots D$ are rank-$r_{d}$ matrices, by incorporating the covering number of low-rank matrices into (60) in Appendix E, we can derive
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
[…]
where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]\ \text{are unknown with }\text{rank}(\bm{B}_{d})=r_{d},d=0,\dots D,\ \text{and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$. Leveraging low-rank structures offers a potential avenue for improving error performance, as the recovery error may be reduced below (23) when the ranks $r_{d},d=0,\dots,D$ are small.
$(ii)$ Second, when $\mathcal{S}_{i},i\in[D]$, are unknown, introducing the covering number of $\{\mathcal{S}_{i}\}_{i=1}^{D}$ into the proof in Appendix E, similar to (23), we can arrive at
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
[…]
where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{N_{D+1}\times K\times N_{0}}:\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d=0,\dots,D,\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are unknown}\}$. In this analysis, we assume that all IRSs are equipped with active elements. However, in practice, when adjustments are required for only a subset of IRSs, denoted as $\mathcal{S}_{i}$ for $i\in\Omega$, where $\Omega\subseteq\{1,\dots,D\}$, the recovery error can be further reduced. Specifically, the term $\sum_{d=1}^{D}KN_{d}$ is replaced by $\sum_{i\in\Omega}KN_{i}$.
Discussion
Beyond the IRS-assisted MIMO system, the tensor-based ToT model can also be effectively applied to a variety of typical communication systems, including direct-sequence code-division multiple-access systems [72], space-time frequency MIMO systems [73], and cooperative/relay systems [74, 75]. Compared with traditional representations, tensor-based methods offer a more compact and efficient form for modeling and processing. Furthermore, by following the same analytical approach above, the optimal recovery error for channel estimation in these systems can be derived, as it is proportionate to the degrees of freedom of the channel matrices, assuming the RIP condition is satisfied using certain transmitted pilot signals. These results offer a theoretical guarantee for robustness and potential accuracy in estimating channel characteristics.
IV Nonconvex Optimization Algorithms
In this section, our focus will be on designing algorithms for estimating the channel matrices within the framework of the TT-based ToT regression model. Initially, we will concentrate on the single-hop IRS-assisted MIMO system and subsequently extend our approach to the multi-hop IRS-assisted MIMO system. Note that, given that the following algorithms can be readily extended to handle the unknown $\mathcal{S}$ scenario, our attention is directed towards the known $\mathcal{S}$ case at the receiver.
Single-hop IRS-assisted MIMO system
$(i)$ One common assumption is that the pilot signal matrix $\bm{X}$ is a truncated discrete Fourier transform (DFT) matrix or a semi-unitary matrix, as discussed in previous works [13, 31]. However, we can extend this assumption by considering the existence of the pseudoinverse for $\bm{X}$. Expanding upon the equivalent form of the TT format $\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]$ [76], denoted as $\mathcal{B}^{\langle 1\rangle}=\bm{H}R(\mathcal{S})(\bm{G}\otimes{\bf I}_{K})$ and $\mathcal{B}^{\langle 2\rangle}=({\bf I}_{K}\otimes\bm{H})L(\mathcal{S})\bm{G}$ (The expressions for $L(\mathcal{S})$ and $R(\mathcal{S})$ are provided in (31) and (32) of Appendix A.)  [^6] Let $\mathcal{X}^{\langle i\rangle}\in\mathbb{C}^{(d_{1}\cdots d_{i})\times(d_{i+1}\cdots d_{N})}$ denote the $i$-th unfolding matrix of the tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times\dots\times d_{N}}$. The $(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})$-th element of $\mathcal{X}^{\langle i\rangle}$ is given by $\mathcal{X}^{\langle i\rangle}(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})=\mathcal{X}(s_{1},\dots,s_{N})$. Here, $s_{1}\cdots s_{i}$ and $s_{i+1}\cdots s_{N}$ represent the row and column indices, respectively, where the row index is calculated as $s_{1}+d_{1}(s_{2}-1)+\cdots+d_{1}\cdots d_{i-1}(s_{i}-1)$ and the column index as $s_{i+1}+d_{i+1}(s_{i+2}-1)+\cdots+d_{i+1}\cdots d_{N-1}(s_{N}-1)$., we can formulate the following loss function:
$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}}\|\widehat{\bm{H}}R(\mathcal{S})(\widehat{\bm{G}}\otimes{\bf I}_{K})-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}\|_{F}^{2},$ (25)
$\min_{\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}\|({\bf I}_{K}\otimes\widehat{\bm{H}})L(\mathcal{S})\widehat{\bm{G}}-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}\|_{F}^{2}.$ (26)
Specifically, we apply the following alternating least squares (ALS) algorithm:
$\bm{H}^{(l+1)}=(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}(R(\mathcal{S})(\bm{G}^{(l)}\otimes{\bf I}_{K}))^{\dagger},$
$\bm{G}^{(l+1)}=(({\bf I}_{K}\otimes\bm{H}^{(l+1)})L(\mathcal{S}))^{\dagger}(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}.$
To guarantee the existence of all pseudoinverse matrices, it is imperative that the conditions $N\leq\min\{UMK,LPK\}$ and $T\geq UM$ are met but we note that $N\leq\min\{UMK,LPK\}$ is a sufficient condition, not a necessary one. This is because Equation 12 holds for all coefficients when $T\geq\Omega(UM)$. Additionally, when $\bm{H}$ and $\bm{G}$ exhibit low-rank characteristics, we can employ SVD to truncate small singular values, facilitating the use of low-rank projection in each iteration.
$(ii)$ To address the numerical stability challenges of the pseudoinverse and to accommodate a broader class of pilot signal matrices $\bm{X}$, we can alternatively optimize the following loss function directly:
$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}\atop\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}f(\widehat{\bm{H}},\widehat{\bm{G}})=\frac{1}{T}\|\mathcal{X}([\widehat{\bm{H}},\mathcal{S},\widehat{\bm{G}}])-\mathcal{Y}\|_{F}^{2}.$ (27)
As for (27), we solve the above optimization problem by the following alternating gradient descent (AGD):
$\bm{H}^{(l+1)}=\bm{H}^{(l)}-\mu\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)}),$
$\bm{G}^{(l+1)}=\bm{G}^{(l)}-\mu\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)}),$
where $\mu$ is a step size and we define the Wirtinger gradients $\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)})$ and $\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)})$ as follows:
[…]
×$\times_{2,3}^{2,3}(\mathcal{S}^{*}\times_{3}^{1}(\bm{G}^{(l)})^{*}),$
[…]
×$\times_{1,2}^{1,2}\big{(}(\mathcal{B}^{(l)}\times_{3}^{1}\bm{X}-\mathcal{Y})\times_{3}^{2}\bm{X}^{*}\big{)}.$
In comparison to the ALS algorithm, the AGD does not require additional coefficient constraints, aside from the condition $T\geq\Omega(UM)$ for satisfying the RIP condition. Additionally, it avoids the numerical stability issues associated with the pseudoinverse.
Figure 1: Recovery performance for channel matrices (a) for different $T$ with $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (b) for different $UM$ with $T=100$, $LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (c) for different $LP$ with $T=20$, $UM=K=N=10$ and $\gamma^{2}=10^{-6}$.
Figure 2: Recovery performance for channel matrices (a) for different $N$ with $T=20$, $UM=LP=K=10$ and $\gamma^{2}=10^{-6}$, (b) for different $K$ with $T=20$, $UM=LP=N=10$ and $\gamma^{2}=10^{-6}$, $\mu=0.02$ for the $1$ IRS scenario and $\mu=0.1$ for the $2$ IRSs scenario, (c) for different $\gamma^{2}$ with $T=20$ and $UM=LP=K=N=10$.
Multi-hop IRS-assisted MIMO system
In a multi-hop IRS-assisted MIMO system, where channel matrices and IRSs are represented in (20), a closed-form solution for the ALS method is not feasible. Consequently, we primarily consider the AGD method and the following loss function:
$\operatorname*{\text{arg~{}min}}_{\widehat{\bm{B}}_{D}\in\mathbb{C}^{LP\times N_{D}},\widehat{\bm{B}}_{0}\in\mathbb{C}^{N_{1}\times UM},\atop\widehat{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]}$ $g(\widehat{\bm{B}}_{D},\dots,\widehat{\bm{B}}_{0})$ (28)
[…]
We solve this nonconvex optimization problem by the following AGD algorithm for $d=0,\dots,D$:
$\bm{B}_{d}^{(l+1)}=\bm{B}_{d}^{(l)}\!\!-\!\mu\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D}^{(l+1)},\dots,\bm{B}_{d+1}^{(l+1)},\bm{B}_{d}^{(l)},\dots,\bm{B}_{0}^{(l)}),$
where $\mu$ is a step size and the Wirtinger gradient $\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\\ \dots,\bm{B}_{0})$ is defined as:
$\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})=\frac{1}{T}\sum_{k=1}^{K}\mathcal{S}_{d+1}^{\rm H}(:,k,:)\bm{B}_{d+1}^{\rm H}\cdots$
$\hskip 14.22636pt\mathcal{S}_{D}^{\rm H}(:,k,:)\bm{B}_{D}^{\rm H}\bigg{(}\mathcal{X}([\bm{B}_{D},\mathcal{S}_{D}(:,k,:),\dots,\bm{B}_{1},\mathcal{S}_{1}(:,k,:),$
$\hskip 14.22636pt\bm{B}_{0}])-\widetilde{\mathcal{Y}}(:,k,:)\bigg{)}\bm{X}^{\rm H}\bm{B}_{0}^{\rm H}\cdots\bm{B}_{d-1}^{\rm H}\mathcal{S}_{d}^{\rm H}(:,k,:).$
In line with the single-hop IRS-assisted MIMO system, AGD in this scenario similarly operates without coefficient constraints, except for the condition $T\geq\Omega(UM)$ for satisfying the RIP condition.
We note that for low-rank channel estimation, iterative hard thresholding combined with truncated singular value decomposition [77] can be effectively integrated into the AGD method.
V Experimental Results
In this section, we conduct numerical experiments to evaluate the effectiveness of Theorem 2 and Theorem 5. We consider $1$ IRS and $2$ IRSs assisted MIMO systems, where each element of channel matrices $\bm{H}$, $\bm{G}$, $\bm{B}_{0}$, $\bm{B}_{1}$, $\bm{B}_{2}$, and pilot signal matrix $\bm{X}$ follows independent and identically complex standard normal distribution. Subsequently, we normalize the above channel matrices to have a unit Frobenius norm. To effectively estimate the channel matrices in the $1$ IRS assisted MIMO system, we apply the ALS, AGD, tensor train singular value decomposition (TT-SVD) [37], and tensor train cross-approximation (TT CA) via the greedy restricted cross-interpolation (GRCI) algorithm [78]. As two representative methods for TT decomposition, TT-SVD and TT-CA are employed for the tensor factorization of $\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger}$. In the case of the $2$ IRSs model, we employ the AGD algorithm. We initialize ALS and AGD with random matrices, with elements generated from a complex standard normal distribution. To simplify parameter selection and notation in the $2$ IRSs system, we set $N_{1}=N_{2}=N$ and denote $\mathcal{B}_{2}=\mathcal{B}$, $\mathcal{B}_{2}^{\star}=\mathcal{B}^{\star}$. The number of iterations for the ALS, AGD ($1$ IRS), and AGD ($2$ IRSs) algorithms are set to $100$, $10^{5}$, and $2\times 10^{5}$, respectively. For the AGD algorithm, we set $\mu=0.2$ for the $1$ IRS scenario and $\mu=1$ for the $2$ IRSs scenario. For each experimental setting, we conduct 20 Monte Carlo trials and then take the average over the 20 trials to report the results.
In the first experiment, we compare the recovery performance of different methods with various values of $T$, $UM$, and $LP$ in Figure 1. These plots illustrate that the recovery error tends to decrease with larger values of $T$, or smaller values of $UM$ and $LP$, consistent with our theoretical findings presented in Theorem 2 and Theorem 5. Additionally, we observe the following: $(i)$ when $T$ approaches $UM$, the recovery error could worsen since RIP conditions (9) and (21) are not satisfied; $(ii)$ AGD (1 IRS) and ALS exhibit similar recovery performance except for when $T=UM$; $(iii)$ the recovery error of AGD (2 IRSs) is larger than AGD (1 IRS) due to the introduction of an additional channel matrix, leading to more error as shown in Theorem 5; $(iv)$ compared to the ALS and AGD algorithms, TT-SVD and TT CA demonstrate poorer recovery performance. This is because TT-SVD and TT CA do not utilize the known information of $\mathcal{S}$ or $\mathcal{S}_{i},i=1,2$, resulting in a larger error. Furthermore, the recovery error of TT-CA is larger than that of TT-SVD, as TT-SVD utilizes the full information of the channel tensor, whereas TT-CA does not [79]; $(v)$ from Figure 1, it is evident that successful recovery is not dependent on $LP$, as even with $LP=100$, the recovery remains stable when $T=20$.
In the second experiment, we examine the recovery performance of different methods across varying values of $N$ and $K$. Keeping a fixed number of time slots $T=20$, we observe stable recovery errors with increasing $N$ and $K$. However, it is noteworthy that as $K$ increases, the recovery performance generally improves, except for TT-SVD and TT CA. This is attributed to the fact that the coherence time $T_{s}=KT$ becomes larger as $K$ increases while $T$ remains fixed, thereby enhancing performance. However, factorization approaches like TT-SVD and TT CA struggle to effectively denoise as $K$ increases. This is due to the possibility that the estimated $\widehat{\mathcal{S}}$ or $\widehat{\mathcal{S}}_{1},\widehat{\mathcal{S}}_{2}$ may introduce more noise under these conditions. This highlights a limitation of tensor factorization in recovering channel matrices of IRS-based systems.
In the third experiment, we compare the recovery performance of different methods across various noise variances $\gamma^{2}$. In Figure 2, we observe that, except for TT CA, the recovery error exhibits a linear relationship with $\gamma^{2}$, consistent with the analysis in Theorem 2 and Theorem 5. Ultimately, while the ALS achieves performance similar to the AGD, the AGD may offer greater stability, especially in time-varying systems where the pseudoinverse of the input matrix may not be well-defined.
Figure 3: Recovery performance for channel matrices for different $D$ with $K=N_{0}=\cdots N_{D}=10$, $\gamma^{2}=10^{-6}$ and $\mu$ sequentially selected from the set $\{0.2,1,5,10,15\}$.
In the final experiment, we compare the recovery performance of AGD for different numbers of IRSs $D$. As shown in Figure 3, the recovery error increases linearly with $D$, aligning with the theoretical analysis in Theorem 5. Additionally, as $T$ increases, the recovery error decreases. Thus, for fixed $T$ time slots, a trade-off emerges between mitigating propagation attenuations with multiple IRSs and minimizing recovery error.
VI Conclusion
This paper addresses the significant gap in theoretical error analysis concerning channel estimation for IRS-assisted MIMO systems. By establishing the equivalence between these systems and tensor train-based tensor-on-tensor (ToT) regression, we provide insights into the fundamental factors crucial for the successful recovery of channel matrices. Our analysis highlights the pivotal role of the relationship between the number of user terminals or base stations and the number of time slots for ensuring stable recovery. Furthermore, we extend our investigation to consider low-rank channel matrices and unknown IRS, enhancing the applicability of our findings. Additionally, through our exploration of a multi-hop IRS scheme, we evaluate corresponding recovery errors, shedding light on the performance of such configurations. Finally, to validate our theoretical conclusions, we propose and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–demonstrating their effectiveness through experimental results.
An important area for future research involves conducting error analysis on two-way, two-hop IRS-assisted MIMO systems [80, 81, 82]. Unlike the one-way multi-hop IRS-assisted model discussed in this paper, the optimization problem in the two-way scenario, where channels between base stations/user terminals and IRSs are accessible, consists of two ToT regression problems. Thus, our current analytical approach cannot be directly extended to this setup. Furthermore, another challenging aspect for future investigation is the local convergence analysis and global geometry of the IRS-assisted optimization problem. Despite the promising experimental results obtained even with random initialization, unlike the rotation ambiguity commonly encountered in standard nonconvex matrix/tensor recovery [45, 46, 47, 48], a diagonal (scaling) ambiguity exists between channel matrices, as noted in [36, Section 2]. This ambiguity poses a challenge in analyzing the optimization problem at hand.
Appendix A Introduction of the TT decomposition
For an order-$D$ tensor $\mathcal{B}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$, the $(s_{1},\dots,s_{D})$-th element of $\mathcal{B}$ in the TT format can be expressed as the following matrix product form [37]
$\mathcal{B}(s_{1},\dots,s_{D})=\bm{B}_{1}(:,s_{1},:)\bm{B}_{2}(:,s_{2},:)\cdots\bm{B}_{D}(:,s_{D},:),$ (29)
where tensor factors ${\bm{B}}_{d}\in\mathbb{C}^{r_{d-1}\times N_{d}\times r_{d}},d=1,\dots,D$ with $r_{0}=r_{D}=1$. Thus, the TT format can be represented by $D$ tensor factors $\{{\bm{B}}_{d}\}_{d\geq 1}$, with a total of $O(D\overline{N}\overline{r}^{2})$ parameters, where $\overline{N}=\max_{d}N_{d}$ and $\overline{r}=\max_{d}r_{d}$. In addition, for any two TT format tensors ~,^$\widetilde{\mathcal{B}},\widehat{\mathcal{B}}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ with factors $\{\widetilde{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widetilde{r}_{d-1}\times\widetilde{r}_{d}}\}$ and $\{\widehat{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widehat{r}_{d-1}\times\widehat{r}_{d}}\}$, each element of the summation $\mathcal{B}=\widetilde{\mathcal{B}}+\widehat{\mathcal{B}}$ can be represented by
$\mathcal{B}(s_{1},\dots,s_{D})=\begin{bmatrix}\widetilde{\bm{B}}_{1}(s_{1})\!\!\!\!&\widehat{\bm{B}}_{1}(s_{1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{2}(s_{2})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{2}(s_{2})\end{bmatrix}$
$\cdots\begin{bmatrix}\widetilde{\bm{B}}_{D-1}(s_{D-1})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{D-1}(s_{D-1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{D}(s_{D})\\ \widehat{\bm{B}}_{D}(s_{D})\end{bmatrix},$ (30)
which implies that $\mathcal{B}$ can also be represented in the TT format with ranks $r_{d}\leq\widetilde{r}_{d}+\widehat{r}_{d}$ for $d=1,\dots,D-1$.
Since any tensor can be decomposed in the TT format (29) with sufficiently large TT ranks [37, Theorem 2.1], the decomposition of a tensor $\mathcal{B}$ into the form (29) is generally not unique: not only are the factors ${\bm{B}_{d}(:,s_{d},:)}$ not unique, but also the dimensions of these factors can vary. To introduce the factorization with the smallest possible dimensions $\bm{r}=(r_{1},\ldots,r_{D-1})$, for convenience, for each $d$, we put $\{{\bm{B}_{d}(:,s_{d},:)}\}_{s_{d}=1}^{N_{d}}$ together into the following two forms
$L(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)\\ \vdots\\ \bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{(r_{d-1}N_{d})\times r_{d}},$ (31)
$R(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)&\cdots&\bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{r_{d-1}\times(N_{d}r_{d})},$ (32)
where $L(\bm{B}_{d})$ and $R(\bm{B}_{d})$ are often called the left and right unfoldings of $\bm{B}_{d}$, respectively, if we view $\bm{B}_{d}$ as a tensor. We say the decomposition (29) is minimal if the rank of the left unfolding matrix $L(\bm{B}_{d})$ is $r_{d}$ and the rank of the right unfolding matrix $R(\bm{B}_{d})$ is $r_{d-1}$ for all $d$. The dimensions $\bm{r}=(r_{1},\dots,r_{D-1})$ of such a minimal decomposition are called the TT ranks of $\mathcal{B}$. According to [50], there is exactly one set of ranks $\bm{r}$ that $\mathcal{B}$ admits a minimal TT decomposition.
Appendix B Proof of Theorem 1
Proof.
By substituting the concentration inequality from [64, Theorem 2.3] with the concentration inequality for complex-valued subgaussian random variables, when (9) holds true, we can assert with a probability of $1-e^{-cT}$ that:
[…]
Due to $\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\sum_{p,k}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ and $\|\mathcal{B}\|_{F}^{2}=\sum_{p,k}\|\mathcal{B}(p,k,:)\|_{F}^{2}$, this completes the proof. ∎
Appendix C Proof of Theorem 2
Proof.
Using (7), we have
[…]
$-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})-\mathcal{X}(\mathcal{B}^{\star})-\mathcal{W}\|_{F}^{2}$
[…]
$+\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})\|_{F}^{2}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})\|_{F}^{2}$
[…]
which further implies that
$\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2}\leq\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|.$ (35)
According to (11), we can directly obtain
$\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2}\geq(1-\delta_{UM})\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}^{2}.$ (36)
For the right-hand side of (35), a straightforward application of the Cauchy-Schwarz inequality $|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|\leq\|\mathcal{W}\|_{F}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}$ is not adequate to fully elucidate the interplay of all parameters. We address this issue by using the covering argument to bound $|\langle\mathcal{W},\mathcal{X}(\mathcal{B}-\mathcal{B}^{\star})\rangle|$ for all possible $\mathcal{B}$.
Now, we rewrite $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$ as following:
$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$ (37)
[…]
where $\bm{X}^{*}$ is the conjugate matrix of $\bm{X}$ and $\mathcal{W}\times_{3}^{2}\bm{X}^{*}=\sum_{t=1}^{T}\mathcal{W}(:,:,t)\bm{X}^{*}(:,t)$. According to [48, eq.(44)], the last line follows $\|\mathcal{H}\|_{F}=\|\bm{H}_{2}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
To begin, according to [83], we can construct an $\epsilon$-net $\{\bm{H}_{1}^{(1)},\dots,\bm{H}_{1}^{(n_{1})}\}$ with the covering number $n_{1}\leq(\frac{4+\epsilon}{\epsilon})^{LPN}$ for the set of factors $\{\bm{H}_{1}\in\mathbb{C}^{LP\times N}:\|\bm{H}_{1}\|\leq 1\}$ such that
$\sup_{\bm{H}_{1}:\|\bm{H}_{1}\|\leq 1}\min_{p_{1}\leq n_{1}}\|\bm{H}_{1}-\bm{H}_{1}^{(p_{1})}\|\leq\epsilon.$ (38)
Similarly, we can construct $\epsilon$-net $\{\bm{H}_{2}^{(1)},\dots,\bm{H}_{2}^{(n_{2})}\}$ with the covering number $n_{2}\leq(\frac{2+\epsilon}{\epsilon})^{UMN}$ for $\{\bm{H}_{2}\in\mathbb{C}^{N\times UM}:\|\bm{H}_{2}\|_{F}\leq 1\}$ such that
$\sup_{\bm{H}_{2}:\|\bm{H}_{2}\|_{F}\leq 1}\min_{p_{2}\leq n_{2}}\|\bm{H}_{2}-\bm{H}_{2}^{(p_{2})})\|_{F}\leq\epsilon.$ (39)
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number
$n_{1}n_{2}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{LPN+UMN}$ (40)
for any TT format tensor $\mathcal{H}=[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}]\in\mathbb{B}_{2N,\widetilde{\mathcal{S}}}$.
Denote by $A$ the value of (37), i.e.,
$[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]=\operatorname*{\text{arg~{}max}}_{\|\bm{H}_{1}\|\leq 1,\|\bm{H}_{2}\|_{F}\leq 1,\atop\|L(\widetilde{\mathcal{S}})\|\leq 1}1/T|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$ (41)
$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (42)
Using $\mathcal{I}$ to denote the index set $[n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
$\|\widetilde{\bm{H}}_{1}-\bm{H}_{1}^{(p_{1})}\|\leq\epsilon,\ \ \ \text{and}\ \ \ \|\widetilde{\bm{H}}_{2}-\bm{H}_{2}^{(p_{2})}\|_{F}\leq\epsilon,$ (43)
and taking $\epsilon=\frac{1}{4}$ gives
[…]
$+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]-[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
[…]
$+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1}-\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]$
$+[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}-\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
[…]
Note that each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$. When conditional on $\bm{X}$, for any fixed $\mathcal{H}^{(p)}=[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}]\in\mathbb{C}^{LP\times K\times UM}$, $\frac{1}{T}\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle$ has complex normal distribution with zero mean and variance $\frac{\gamma^{2}\|\mathcal{X}(\mathcal{H}^{(p)})\|_{F}^{2}}{T^{2}}$, which implies that
$\operatorname{\mathbb{P}}\left(\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq t|\bm{X}\right)\leq e^{-\frac{T^{2}t^{2}}{2\gamma^{2}\|\mathcal{X}(\mathcal{H}^{(p)})\|_{F}^{2}}}.$ (45)
[…]
$\operatorname{\mathbb{P}}\left(\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq t|F\right)\leq e^{-\frac{Tt^{2}}{2(1+\delta_{UM})\gamma^{2}}}.$ (46)
We now apply this tail bound to (44) and get
[…]
where $c_{1}$ is a constant and based on the assumption $\epsilon=\frac{1}{4}$ in (44), $\frac{4+\epsilon}{\epsilon}=17$.
Hence, we can take $t=\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma$ with a constant $c_{2}$ and further derive
$\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma\right)$ (48)
[…]
where $c_{i},i=3,4,5$ are constants. Note that $P(F)$ follows Theorem 1.
Combing (36), we can obtain
^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)}.$ (49)
Appendix D Proof of Theorem 3
Proof.
Suppose we can find a set of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that $\min_{j\neq k}\|\mathcal{B}^{j}-\mathcal{B}^{k}\|_{F}\geq s$. According to [84, Theorem 4], when each element of $\bm{X}$ and $\mathcal{W}$ respectively follow $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, we have
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}$ (50)
[…]
Next, we consider one construction for the sets of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that we can obtain a proper lower bound for $\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$ and a proper upper bound for $\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$. Given the equivalence between any two tensors $\mathcal{B}^{j_{1}}=[\bm{A}_{1}^{j_{1}},\mathcal{S},\bm{A}_{2}]$, $\mathcal{B}^{j_{2}}=[\bm{A}_{1}^{j_{2}},\mathcal{S},\bm{A}_{2}]$ and their canonical forms [48], the TT-SVD can be applied to transform these tensors into canonical formats $\mathcal{B}^{j_{1}}=[\widetilde{\bm{A}}_{1}^{j_{1}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ and $\mathcal{B}^{j_{2}}=[\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ with $R(\widetilde{\mathcal{S}})R^{\rm H}(\widetilde{\mathcal{S}})={\bf I}_{N}$ and $\widetilde{\bm{A}}_{2}\widetilde{\bm{A}}_{2}^{\rm H}={\bf I}_{N}$. Given any $\delta>0$ and $N\geq C^{\prime}$ by [85, Lemma 5], we can construct a set of $\{\widetilde{\bm{A}}_{1}^{1},\dots,\widetilde{\bm{A}}_{1}^{n_{0}}\}$ with cardinality $n_{0}\geq\frac{1}{4}e^{\frac{LPN}{128}}$ such that: (1) $\|\widetilde{\bm{A}}_{1}^{j}\|_{F}=\delta$ holds for all $j=1,\dots,n_{0}$, (2) $\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta$ for all $j_{1},j_{2}\in[n_{0}],j_{1}\neq j_{2}$.
Due to $\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}=\|[\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]\|_{F}=\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}$, we can further get
$\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}=\max_{j_{1}\neq j_{2}}\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\leq 2\delta,$
$\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}=\min_{j_{1}\neq j_{2}}\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta.$ (51)
Then we plug (D) into (50) and have
[…]
where the second inequality with a constant $c_{2}>0$ follows $\delta=c_{3}\sqrt{\frac{LPN}{m}}\gamma$ for a constant $c_{3}>0$.
Similarly, we can utilize the TT-SVD to transform $\mathcal{B}^{j_{1}}=[\bm{A}_{1}^{j_{1}},\mathcal{S},\bm{A}_{2}]$ and $\mathcal{B}^{j_{2}}=[\bm{A}_{1}^{j_{2}},\mathcal{S},\bm{A}_{2}]$ into orthonormal formats $\mathcal{B}^{j_{1}}=[\widetilde{\bm{A}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}^{j_{1}}]$ and $\mathcal{B}^{j_{2}}=[\widetilde{\bm{A}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}^{j_{2}}]$ with $L^{\rm H}(\widetilde{\mathcal{S}})L(\widetilde{\mathcal{S}})={\bf I}_{N}$ and $\widetilde{\bm{A}}_{1}^{\rm H}\widetilde{\bm{A}}_{1}={\bf I}_{N}$. Then following the previous analysis, we can get
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{UMN}{T}}\gamma\bigg{)}.$ (53)
Finally, we sum (52)-(53) and calculate the average to derive
$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{2T}}\gamma\bigg{)}.$ (54)
Appendix E Proof of Theorem 5
Proof.
Utilizing the identical derivation as (35), we can directly acquire
$\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\|_{F}^{2}\leq\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|.$ (55)
First, according to Theorem 4, we have
$\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\|_{F}^{2}\geq(1-\delta_{UM})\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}^{2}.$ (56)
Additionally, we can also rewrite $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$ as follows:
$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$
$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N}^{D},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (57)
The last line follows that $\|\mathcal{H}\|_{F}^{2}=\|\bm{H}_{0}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
Next, we can construct an $\epsilon$-net $\{\bm{H}_{i}^{(1)},\dots,\bm{H}_{i}^{(n_{i})}\}$ with the covering number $n_{i}\leq(\frac{4+\epsilon}{\epsilon})^{N_{i}N_{i+1}},i\in[D]$ for the set of factors $\{\bm{H}_{i}\in\mathbb{C}^{N_{i+1}\times N_{i}}:\|\bm{H}_{i}\|\leq 1\}$ such that
$\sup_{\bm{H}_{i}:\|\bm{H}_{i}\|\leq 1}\min_{p_{i}\leq n_{i}}\|\bm{H}_{i}-\bm{H}_{i}^{(p_{i})}\|\leq\epsilon.$ (58)
In addition, we can construct an $\epsilon$-net $\{\bm{H}_{0}^{(1)},\dots,\bm{H}_{0}^{(n_{0})}\}$ with the covering number $n_{0}\leq(\frac{2+\epsilon}{\epsilon})^{N_{0}N_{1}}$ for the set of factors $\{\bm{H}_{0}\in\mathbb{C}^{N_{1}\times N_{0}}:\|\bm{H}_{0}\|_{F}\leq 1\}$ such that
$\sup_{\bm{H}_{0}:\|\bm{H}_{0}\|_{F}\leq 1}\min_{p_{0}\leq n_{0}}\|\bm{H}_{0}-\bm{H}_{0}^{(p_{0})}\|_{F}\leq\epsilon.$ (59)
Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{0}\cdots n_{D})}\}$ with covering number
$n_{0}\cdots n_{D}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{\sum_{d=0}^{D}N_{d}N_{d+1}}$ (60)
for any $\mathcal{H}\in\mathbb{B}_{2N,\{\widetilde{\mathcal{S}}_{i}\}}^{D}$.
Denote by $A$ the value of (E), i.e.,
$[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]$ (62)
[…]
$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$
Using $\mathcal{I}$ to denote the index set $[n_{0}]\times\cdots\times[n_{D}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{0},\dots,p_{D})\in\mathcal{I}$ such that
$\|\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})}\|\leq\epsilon,i\in[D]\ \text{and}\ \|\widetilde{\bm{H}}_{0}-\bm{H}_{0}^{(p_{0})}\|_{F}\leq\epsilon,$ (63)
and taking $\epsilon=\frac{1}{2(D+1)}$ gives
[…]
$+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]$
$-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
[…]
$+\frac{1}{T}\bigg{|}\sum_{i=0}^{D}\bigg{\langle}[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\bm{H}_{D-1}^{(p_{D-1})},\dots,\bm{H}_{i+1}^{(p_{D_{i+1}})},\widetilde{\mathcal{S}}_{i+1},$
$\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})},\widetilde{\mathcal{S}}_{i},\widetilde{\bm{H}}_{i-1},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\bigg{\rangle}\bigg{|}$
[…]
$+(D+1)\epsilon A$
[…]
where the first equation follows the fact that $(l,k,u)$-th element of $[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}]$ can be decomposed into $\sum_{i=0}^{D}\bm{H}_{D}^{(p_{D})}(l,:)\widetilde{\mathcal{S}}_{D}(:,k,:)\bm{H}_{D-1}^{(p_{D-1})}\\ \cdots\bm{H}_{i+1}^{(p_{D_{i+1}})}\widetilde{\mathcal{S}}_{i+1}(:,k,:)(\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})})\widetilde{\mathcal{S}}_{i}(:,k,:)\widetilde{\bm{H}}_{i-1}\widetilde{\mathcal{S}}_{1}(:,k,:)\widetilde{\bm{H}}_{0}(:,u)$.
Based on the same analysis of (48), we can obtain
$\operatorname{\mathbb{P}}\left(A\leq\frac{c_{1}\sqrt{(1+\delta_{UM})(\sum_{d=1}^{D-1}N_{d}N_{d+1})\log D}}{\sqrt{T}}\gamma\right)$ (65)
[…]
where $c_{i},i=1,2$ are constants.
Combing (56), we arrive at
^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$ (66)
[…]

### v2

@@
 As intelligent reflecting surface (IRS) has emerged as a new and promising technology capable of configuring the wireless environment favorably, channel estimation for IRS-assisted multiple-input multiple-output (MIMO) systems has garnered extensive attention in recent years.
-While various algorithms have been proposed to address this challenge, there is a lack of rigorous theoretical error analysis.
+Despite the development of numerous algorithms to address this challenge, a comprehensive theoretical characterization of the optimal recovery error is still lacking.
 This paper aims to address this gap by providing theoretical guarantees in terms of stable recovery of channel matrices for noisy measurements.
-We begin by establishing the equivalence between IRS-assisted MIMO systems and a compact tensor train (TT)-based tensor-on-tensor (ToT) regression.
+We begin by establishing the equivalence between IRS-assisted MIMO systems in the uplink scenario and a compact tensor train (TT)-based tensor-on-tensor (ToT) regression.
 Building on this equivalence, we then investigate the restricted isometry property (RIP) for complex-valued subgaussian measurements.
-Our analysis reveals that successful recovery hinges on the relationship between the number of user terminals (in the uplink scenario) or base stations (in the downlink scenario) and the number of time slots during which channel matrices remain invariant.
-Utilizing the RIP condition, we analyze the theoretical recovery error for the solution to a constrained least-squares optimization problem, including upper error bound and minimax lower bound, demonstrating that the error decreases inversely with the number of time slots and increases proportionally with the number of unknown elements in the channel matrices.
-In addition, we extend our error analysis to two more specialized IRS-assisted MIMO systems, incorporating low-rank channel matrices or an unknown IRS.
-Furthermore, we explore a multi-hop IRS scheme and analyze the corresponding recovery errors.
-Finally, we introduce and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–to validate our conclusions through simulations.
-Index Terms:
-Intelligent reflecting surface, restricted isometry property, error analysis, tensor train-based tensor-on-tensor regression.
+Our analysis reveals that successful recovery hinges on the relationship between the number of user terminals and the number of time slots during which channel matrices remain invariant.
+Utilizing the RIP condition, we establish a theoretical upper bound on the recovery error for solutions to the constrained least-squares optimization problem, as well as a minimax lower bound for the considered model.
+Our analysis demonstrates that the recovery error decreases inversely with the number of time slots, and increases proportionally with the total number of unknown entries in the channel matrices, thereby quantifying the fundamental trade-offs in channel estimation accuracy.
+In addition, we explore a multi-hop IRS scheme and analyze the corresponding recovery errors.
+Finally, we have performed numerical experiments to support our theoretical findings.
 I Introduction
@@
 One particular hurdle stems from the assumption that an IRS typically comprises passive elements, thereby necessitating the receiver to estimate the cascaded channel based on pilots transmitted by the transmitter through the IRS.
-Recently, numerous matrix-based methods for channel estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] have been proposed and demonstrated commendable performance in practical implementation.
+Recently, numerous matrix-based methods for channel estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] have been proposed and demonstrated commendable performance in practical implementation.
 Building on these advancements, another promising approach leverages tensor models for channel estimation.
@@
 On the other hand, since CP decomposition combined with an identity tensor can be viewed as a special case of the Tucker decomposition [36], [33] constructed a Tucker-based IRS-assisted MIMO model and applied a quasi-optimal method–Higher Order Singular Value Decomposition–to achieve efficient channel estimation.
-Although various optimization algorithms for IRS-assisted MIMO systems have demonstrated excellent performance, existing analyses of stable recovery focus solely on the required number of time slots by exploiting the uniqueness of solutions in matrix-based models [36] or the uniqueness conditions of tensor decompositions [30].
-Despite these advances, there is no comprehensive error analysis for channel estimation in IRS-assisted MIMO systems, leading to the following critical question:
-Question: What is the optimal recovery error for the channel estimation of the IRS-assisted MIMO system?
-Our contribution:
-In this paper, we affirmatively address the main question through error analyses in the tensor-based muliple-access/multi-user IRS-assisted MIMO system.
-Instead of using CP or Tucker decompositions, we first rigorously establish that the combination of channel matrices and IRS is equivalent to the tensor train (TT) format [37], which has been extensively utilized in quantum physics for efficiently representing quantum states [38, 39, 40].
-Additionally, we demonstrate that the IRS-assisted MIMO model aligns with the tensor-on-tensor (ToT) regression model [41, 42, 43] without any assumptions regarding the pilot signals.
-This mode extends the classical matrix/tensor regression framework, in which the response is the vector [44, 45, 46, 47, 48, 49], to accommodate transmitted signals (represented by the second word “tensor” in the name) and the received signals (represented by the second word “tensor” in the name) as tensors.
-Building on the definitions outlined above, the channel estimation in a standard IRS-assisted MIMO model can be equivalently viewed as a recovery problem within the TT-based ToT regression model.
-Importantly, while other matrix/tensor-based models can be equivalently transformed into this framework, the TT-based ToT regression model is particularly advantageous because its canonical TT format [50, Theorem 1] facilitates theoretical analysis and the ToT model provides a compact representation that streamlines the design of optimization methods tailored to various pilot signal configurations.
-Consequently, this paper primarily focuses on analyzing this model.
-Building on the previous discussion, our problem is reformulated as a recovery problem within the TT-based ToT regression model, representing a multiple-access/multi-user MIMO communication system assisted by an IRS in an uplink communication scenario.
-Towards that goal, we initially establish the restricted isometry property (RIP) for complex-valued subgaussian measurements, where each pilot signal element is a complex-valued subgaussian random variable.
+Various optimization algorithms for IRS-assisted MIMO systems have exhibited strong empirical performance.
+However, the theoretical analysis of stable recovery has largely focused on identifying the required number of time slots by leveraging conditions that ensure the uniqueness of the solution [31, 37, 33].
+A lower bound on the estimation error based on the Cramér–Rao bound (CRB) was established in [31].
+However, the lower bound could be suboptimal since it is obtained by analyzing a convex surrogate of the original nonconvex least-squares problem, which alters the intrinsic structure of the estimation problem and, consequently, fails to capture its underlying degrees of freedom (see Section II-D for a detailed discussion).
+To the best of our knowledge, a rigorous optimal error analysis for channel estimation in IRS-assisted MIMO systems is still lacking.
+This raises the following fundamental question:
+Question: What is the optimal recovery error for channel estimation in IRS-assisted MIMO systems, and can this limit be achieved by appropriately formulating the underlying estimation problem?
+I-A Our contribution:
+In this paper, we affirmatively address the main question through error analyses in the muliple-access/multi-user IRS-assisted MIMO system.
+We begin by formally establishing that the combination of the channel matrices and the IRS structure can be naturally represented in the tensor train (TT) format [38].
+Based on this equivalence, we reformulate the IRS-assisted MIMO model as a TT-based tensor-on-tensor (ToT) regression model [39, 40].
+This reformulation simplifies our analysis and naturally leads to a recovery problem under the TT-based ToT regression model, which captures the uplink communication scenario in a multiple-access/multi-user IRS-assisted MIMO system.
+To advance this analysis, we initially establish the restricted isometry property (RIP) for complex-valued subgaussian measurements, where each pilot signal element is a complex-valued subgaussian random variable.
 This analysis reveals an optimal size relationship between user terminals (UTs) and time slots.
-Notably, for a downlink communication scenario, RIP relates to base stations (BSs) and time slots.
-Subsequently, we investigate channel matrices recovery of the least square optimization problem from subgaussian measurements and derive recovery bounds relative to the number of time slots.
-Our findings illustrate that recovery error is inversely proportionate to the number of time slots, and successful recovery depends on the relationship between UTs and time slots.
-Aside from aforementioned constraints, our recovery analyses remain free from other coefficient constraints.
-Moreover, we extend recovery error considerations to include low-rank channel matrices [30] and unknown IRS [51].
+We then investigate the recovery of channel matrices via constrained least squares minimization from subgaussian measurements, deriving optimal recovery error bounds that are proportional to the degrees of freedom of the underlying channels.
+Specifically, we establish an upper bound on the estimation error for constrained least squares objective and further derive a minimax lower bound that depends solely on the underlying IRS-assisted MIMO model, without relying on any specific loss function, thereby further highlighting the optimality of the upper bound.
+Compared to the CRB, the minimax lower bound more accurately reflects the intrinsic nonconvexity of the model, yielding a tight characterization of the recovery error.
+Moreover, we extend recovery error considerations to include low-rank channel matrices [31], sparse matrices [41] and unknown IRS [42].
 Next, we explore a multi-hop IRSs scheme and analyze corresponding recovery errors, laying the groundwork for future multi-hop IRSs communication design.
-Finally, we propose two nonconvex optimization algorithms–alternating least squares and alternating gradient descent and simulations affirm the validity of theoretical analyses.
-Notations
-We use calligraphic letters (e.g., $\mathcal{Y}$) to denote tensors, bold capital letters (e.g., $\bm{Y}$) to denote matrices, except for $\bm{X}_{i}$ which denotes the $i$-th order-$3$ tensor factors in the TT format ($i=2,\dots,N+M-1$), bold lowercase letters (e.g., $\bm{y}$) to denote vectors, and italic letters (e.g., $y$) to denote scalar quantities.
+Finally, simulations affirm the validity of theoretical analyses.
+I-B Notations
+We use calligraphic letters (e.g., $\mathcal{Y}$) to denote tensors, bold capital letters (e.g., $\bm{Y}$) to denote matrices, except for $\bm{X}_{i}$ which denotes the $i$-th order-33 tensor factors in the TT format ($i=2,\dots,N+M-1$), bold lowercase letters (e.g., $\bm{y}$) to denote vectors, and italic letters (e.g., $y$) to denote scalar quantities.
 $a^{*}$ is the complex conjugate of $a$.
@@
 The procedure of $\mathcal{A}\times_{i_{1},\dots,i_{n}}^{j_{1},\dots,j_{n}}\mathcal{B}$ can be viewed as a sequence of $n$ tensor contractions $\mathcal{A}\times_{i_{k}}^{j_{k}}\mathcal{B},k\in[n]$.
-$\|\bm{X}\|$ represents the spectral norm of the matrix $\bm{X}$ For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$.
+The symbol * denotes the Khatri-Rao product, and the operator $\operatorname*{vec}(\cdot)$ refers to the vectorization of a matrix.
+$\|\bm{X}\|$ and $\|\bm{X}\|_{0}$ respectively represent the spectral norm and $\ell_{0}$ norm of the matrix $\bm{X}$.
+For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$.
 For two positive quantities $a,b\in\mathbb{R}$, the inequality $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b=\Omega(a)$ indicates that $b\geq ca$ for some universal constant $c$.
 II Single-Hop IRS-Assisted MIMO System
-Model formulation
-We consider a muliple-access/multi-user MIMO communication system assisted by an IRS, in which the $P$ base stations (BSs) receive the signals transmitted by $U$ user terminals (UTs) via the IRS.
+II-A Model formulation
+We consider a muliple-access/multi-user MIMO communication system assisted by an IRS, in which the $P$ base stations (BSs) receive the signals transmitted by $U$ UTs via the IRS.
 The terminology employed in this paper is based on an uplink communication scenario, where the transmitter is referred to as the UT and the receiver as the BS.
-However, it is important to note that our signal model and theoretical analysis can be equally applied to the downlink case by merely reversing the roles of the transmitter and the receiver.
 Without loss of generality, we assume that each BS and each UT are equipped with the same number of antennas, denoted as $L$ and $M$ respectively.
@@
 $\bm{w}(t)\sim\mathcal{C}\mathcal{N}({\bm{0}},\gamma^{2}{\bf I}_{LP})$ denotes the complex additive white Gaussian noise (AWGN).
-The channel coherence time $T_{s}$ is divided into $K$ blocks, where each block has $T$ time slots so that $T_{s}=KT$ [29].
+The channel training time $T_{s}$ is divided into $K$ blocks, where each block has $T$ time slots so that $T_{s}=KT$ [30].
 Let us define $\bm{y}(k,t)=\bm{y}((k-1)T+t)$ as the received signal at the $t$-th time slot of the $k$-th block, where $t\in[T]$ and $k\in[K]$.
-In addition, we denote $\bm{x}(k,t)$ and $\bm{s}(k,t)$ as the pilot signal and phase shift vectors associated with the $t$-th time slot of the $k$-th block.
-In each block, the user transmits the same pilot signal and the IRS phase shift vector is invariant, but the phase shift vector between blocks is changeable, which means that the IRS has $K$ phase configurations.
-Consequently, we have $\bm{s}(k)=\bm{s}(k,t)$ and $\bm{x}(t)=\bm{x}(k,t)$.
-The received signal is given by
+Similarly, we denote $\bm{x}(k,t)$, $\bm{s}(k,t)$ and $\bm{w}(k,t)$ as the pilot signal, phase shift and noise vectors associated with the $t$-th time slot of the $k$-th block.
+We adopt the structured time-domain training protocol proposed in [30], wherein the IRS phase shift vector remains fixed throughout the TT time slots within each block and changes from block to block.
+Meanwhile, the pilot signals $\begin{bmatrix}\bm{x}(1)\cdots\bm{x}(T)\end{bmatrix}$ are identically repeated across all $K$ blocks.
+Consequently, we have
+$\bm{s}(k)$ $\!\!\!\!=\!\!\!\!$ $\bm{s}(k,t),t\in[T],$
+$\bm{x}(t)$ $\!\!\!\!=\!\!\!\!$ $\bm{x}(k,t),k\in[K].$
+Under these assumptions, the received signal in (1) is given by
 $\bm{y}(k,t)\!=\!\bm{H}\text{diag}(\bm{s}(k))\bm{G}\bm{x}(t)+\bm{w}(k,t),\ t\in[T]\ \text{and}\ k\in[K].$
@@
 Moreover, the CP factorization framework used for single-hop IRS-assisted MIMO systems does not readily extend to multi-hop configurations, presenting additional challenges.
-Compact tensor representation
-To accommodate a more general setting for the pilot signal matrix $\bm{X}$, we can reinterpret (2) as a tensor-on-tensor (ToT) regression model.
+II-B Compact tensor representation
+To accommodate a more general setting for the pilot signal matrix $\bm{X}$ and facilitate the analysis of the optimal recovery error in channel estimation, we can reinterpret (2) as a tensor-on-tensor (ToT) regression model [44, 45, 39, 40] which generalizes the classical matrix/tensor regression framework where the response variable is a vector [46, 47, 48, 49, 50, 51, 52, 53].
 Specifically, by aggregating the received data over $K$ blocks, we can define three tensors as $\mathcal{Y}\in\mathbb{C}^{LP\times K\times T}$, $\mathcal{S}\in\mathbb{C}^{N\times K\times N}$ and $\mathcal{W}\in\mathbb{C}^{LP\times K\times T}$, where $\mathcal{Y}(:,k,:)=\bm{Y}(k)$, $\mathcal{S}(:,k,:)=\text{diag}(\bm{s}(k))$ and $\mathcal{W}(:,k,:)=\bm{W}(k)$.
 Then, the ground-truth channel tensor $\mathcal{B}^{\star}\in\mathbb{C}^{LP\times K\times UM}$ can be defined as follows:
-[…]
-[…]
+$\mathcal{B}^{\star}(p,k,m)$ $\!\!\!\!=\!\!\!\!$ $\bm{H}(p,:)\mathcal{S}(:,k,:)\bm{G}(:,m)$ (3)
+$\!\!\!\!=\!\!\!\!$ $\bm{H}(p,:)\text{diag}(\bm{s}(k))\bm{G}(:,m).$
 Compared to directly analyzing (2), the tensor formulation enables the modeling of correlations in the signal model (2) across different phase configurations indexed by $k$.
@@
 It is evident that the pilot signal (measurement operator) $\bm{X}$ exclusively affects UTs, as the received signal $\mathcal{Y}$ is a tensor of size $LP\times K\times T$ and the channel tensor $\mathcal{B}^{\star}$ in (4) is a tensor of size $LP\times K\times UM$, implying the total number of antennas in the UTs $UM$ should be contingent on the number of time slots $T$.
-Notice that, unlike the construction of CP decomposition, which requires transforming (3) as demonstrated in [29, 30, 52, 13, 31], (3) is inherently represented in the standard tensor train (TT) format as each element of $\mathcal{B}^{\star}$ can be expressed in a matrix product form (a detailed introduction to the TT decomposition is provided in Appendix A).
+Notice that, unlike the construction of CP decomposition, which requires transforming (3) as demonstrated in [30, 31, 43, 13, 32], (3) is inherently represented in the standard tensor train (TT) format [38] as each element of $\mathcal{B}^{\star}$ can be expressed in a matrix product form (a detailed introduction to the TT decomposition is provided in the Appendix A).
 By applying the tensor contraction operation [54], the TT form of $\mathcal{B}^{\star}$ can be expressed as:
@@
 Therefore, we can rewrite (4) as a TT-based ToT regression model.
-[…]
-[…]
-where (×$(\mathcal{B}^{\star}\times_{3}^{1}\bm{X})(p,k,t)=\sum_{j=1}^{UM}\mathcal{B}^{\star}(p,k,j)\bm{X}(j,t)$.
+$\mathcal{Y}=\mathcal{X}(\mathcal{B}^{\star})+\mathcal{W}$ $\!\!\!\!=\!\!\!\!$ $\mathcal{B}^{\star}\times_{3}^{1}\bm{X}+\mathcal{W}$ (6)
+$\!\!\!\!=\!\!\!\!$ $[\bm{H},\mathcal{S},\bm{G}]\times_{3}^{1}\bm{X}+\mathcal{W},$
+where $(\mathcal{B}^{\star}\times_{3}^{1}\bm{X})(p,k,t)=\sum_{j=1}^{UM}\mathcal{B}^{\star}(p,k,j)\bm{X}(j,t)$.
 The linear map $\mathcal{X}(\mathcal{B}^{\star}):\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ models the measurement process.
+Here, we emphasize that the TT-based ToT regression model in (6) serves solely to simplify the theoretical analysis and preserves all the information contained in the original model (2).
 We now consider minimizing the following constrained least squares objective to recover the channel tensor $\mathcal{B}^{\star}$:
-^$\widehat{\mathcal{B}}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\frac{1}{T}\|\mathcal{X}(\mathcal{B})-\mathcal{Y}\|_{F}^{2},$ (7)
+$\widehat{\mathcal{B}}=\operatorname*{\text{arg~min}}_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\frac{1}{T}\|\mathcal{X}(\mathcal{B})-\mathcal{Y}\|_{F}^{2},$ (7)
 where we define a set of order-$3$ TT format tensor as following:
-[…]
+$\mathbb{B}_{N,\mathcal{S}}$ $\!\!\!\!=\!\!\!\!$ $\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\bm{G}\in\mathbb{C}^{N\times UM}\$(8)
 $\text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}.$
-The advantage of the least-squares formulation is that it allows the application of well-established tools from standard matrix/tensor regression models, such as the Restricted Isometry Property (RIP) condition [54, 55, 56, 57, 40], as detailed in the following discussion.
+The advantage of the least-squares formulation is that it allows the application of well-established tools from standard matrix/tensor regression models, such as the RIP condition [55, 56, 57, 58, 50], as detailed in the following discussion.
 Subsequently, we provide a rigorous theoretical analysis, deriving error bounds for channel estimation.
-Error analysis
-One advantage of viewing the channel estimation as the regression model is that analogous to the analysis in standard matrix/tensor regression models, a larger number of time slots $T$ tends to enhance denoising effectiveness since the error bound of channel estimation in the IRS-assisted MIMO model should heavily rely on $T$.
+II-C Optimal error analysis
+One advantage of viewing the channel estimation as the regression model is that analogous to the analysis in standard regression models, a larger number of time slots $T$ tends to enhance denoising effectiveness since the error bound of channel estimation in the IRS-assisted MIMO model should heavily rely on $T$.
 Recently, we analyzed the statistical guarantee of the TT-based ToT regression model in [40].
-However, these results cannot be directly applied to the IRS-assisted MIMO model because recovery errors pertain to the entire tensor recovery.
-Since the phase shift tensor $\mathcal{S}$ is often known, only the channel matrices are affected by noise.
-Consequently, the recovery error should be primarily related to the channel matrices.
-Next, we will establish the upper bound and the minimax lower bound of the recovery error for channel estimation.
+Compared with that model, the IRS-assisted MIMO setting introduces additional structural constraints, as the phase-shift tensor $\mathcal{S}$ is typically known and only the channel matrices are corrupted by noise.
+Consequently, a direct extension of the results in [40] would lead to non-tight recovery bounds that do not properly scale with the degrees of freedom in the unknown channel matrices.
+Since the recovery error is primarily determined by the channel matrices, we next establish both the upper and minimax lower bounds for the corresponding channel estimation problem.
 To facilitate the recovery of channels $\bm{H}$ and $\bm{G}$ from their linear measurements $\mathcal{Y}$, the pilot signal (regression/measurement operator) need adhere to specific properties.
-One such desirable property is the RIP, extensively explored and popularized in compressive regression literature [54, 55, 56, 59].
-This property has also been extended to structured tensors [57, 40].
+One such desirable property is the RIP, extensively explored and popularized in compressive regression literature [55, 56, 57, 59], and later extended to structured tensor models [58, 50].
+From a theoretical perspective, the RIP is typically established by assuming that the measurement operators are drawn randomly from a suitable distribution, with subgaussian measurement ensembles serving as a canonical example.
 Theorem 1.
-Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable  [^2] Note that a complex-valued random variable $X$ is subgaussian if and only if its both real part $\operatorname{Re}\left\{X\right\}$ and imaginary part $\operatorname{Im}\left\{X\right\}$ are real subgaussian random variables.
+A linear operator $\mathcal{X}:\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ is said to satisfy the $UM$-RIP with a constant $\delta_{UM}$ if
+$(1-\delta_{UM})\|\mathcal{B}\|_{F}^{2}\leq\frac{1}{T}\|\mathcal{X}(\mathcal{B})\|_{F}^{2}\leq(1+\delta_{UM})\|\mathcal{B}\|_{F}^{2},$ (9)
+holds for any channel tensor $\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}$. We now consider two cases:
+  • Suppose each element of the pilot signal matrix $\bm{X}$ is an i.i.d.
+complex-valued subgaussian random variable  [^2] Note that a complex-valued random variable $X$ is subgaussian if and only if its both real part $\operatorname{Re}\left\{X\right\}$ and imaginary part $\operatorname{Im}\left\{X\right\}$ are real subgaussian random variables.
 Here are some classical examples of subgaussian distributions.
 (Gaussian) A standard complex Gaussian random variable $X=\operatorname{Re}\left\{X\right\}+i\operatorname{Im}\left\{X\right\}$ with $\operatorname{Re}\left\{X\right\}$ and $\operatorname{Im}\left\{X\right\}$ being independent and following $\mathcal{N}(0,\frac{1}{2})$, is a subgaussian random variable.
-(Bernoulli) A Bernoulli random variable $X$ that takes values $-1$ and $1$ with equal probability is a subgaussian random variable.
+(Bernoulli) A Bernoulli random variable $X$ that takes values 1-1 and $1$ with equal probability is a subgaussian random variable.
 (Digital modulation signals) The phase shift keying (PSK) [60], quadrature amplitude modulation (QAM) [61] and pulse amplitude modulation (PAM) [62] exhibit subgaussian random variable properties.
+Specifically, an i.i.d.
+sequence of PSK, QAM, or PAM symbols can be generated by first drawing a bitstream as an i.i.d.
+random process with zero mean and bounded variance, and then mapping each bit independently to the corresponding constellation points.
+The resulting symbol sequence can thus be treated as an i.i.d.
+complex-valued subgaussian process with a finite subgaussian norm.
 .
-Let $\delta_{UM}\in(0,1)$ be a positive constant.
-Then, for any channel tensor $\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}$, when the number of time slots satisfies
+For a positive constant $\delta_{UM}\in(0,1)$, when the number of time slots satisfies
     $T\geq C\cdot\frac{UM}{\delta_{UM}^{2}},$ (10)
-with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
-$(1-\delta_{UM})\|\mathcal{B}\|_{F}^{2}\leq\frac{1}{T}\|\mathcal{X}(\mathcal{B})\|_{F}^{2}\leq(1+\delta_{UM})\|\mathcal{B}\|_{F}^{2},$ (10)
-where $c$ and $C$ are positive constants.
-The proof is presented in Appendix B.
-Theorem 1 guarantees the RIP for subgaussian measurement ensembles, with the number of time slots $T$ scaling linearly only in relation to the total number of antennas ($UM$) in the UTs  [^3] In a downlink communication scenario, the parameter $T$ will be connected to the total number of antennas ($LP$) in the BSs..
-Subgaussian random variables, including Gaussian, Bernoulli, PSK, QAM, and PAM signals, can be conveniently designed as pilot signals in practical applications.
-Notably, while each element of the Haar-distributed random unitary matrix is a subgaussian random variable, the same does not hold for the elements of the normalized DFT matrix, $\frac{1}{\sqrt{T}}\bm{X}$, a commonly employed pilot signal matrix.
-Nevertheless, the DFT matrix satisfies (10) when $T\geq UM$, rendering the subsequent conclusions with $\delta_{UM}=0$ applicable to them as well.
-Our analysis centers on the subgaussian case to streamline the discussion.
-When RIP holds, then for any two distinct TT format channel tensors $\mathcal{B}_{1},\mathcal{B}_{2}\in\mathbb{B}_{N,\mathcal{S}}$, noting that $\mathcal{B}_{1}-\mathcal{B}_{2}\in\mathbb{B}_{2N,\mathcal{S}}$ is also a TT format channel tensor according to (A) in Appendix A, we have distinct measurements since
-[…]
-[…]
+    with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP with positive constants $c$ and $C$.
+  • Suppose $\frac{1}{\sqrt{T}}\bm{X}$ is the normalized DFT matrix, when
+    $T\geq UM$ (11)
+    is satisfied, we have $\frac{1}{T}\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\|\mathcal{B}\|_{F}^{2}$ with $\delta_{UM}=0$.
+The proof is presented in the Appendix C.
+Theorem 1 guarantees the RIP for subgaussian measurement ensembles and DFT signals, with the number of time slots $T$ scaling linearly only in relation to the total number of antennas ($UM$) in the UTs.
+When RIP holds, then for any two distinct TT format channel tensors $\mathcal{B}_{1},\mathcal{B}_{2}\in\mathbb{B}_{N,\mathcal{S}}$, noting that $\mathcal{B}_{1}-\mathcal{B}_{2}\in\mathbb{B}_{2N,\mathcal{S}}$ is also a TT format channel tensor according to (2) in the Appendix A, we have distinct measurements since
+$\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{1})-\mathcal{X}(\mathcal{B}_{2})\|_{F}^{2}$ $\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{1}-\mathcal{B}_{2})\|_{F}^{2}$ (12)
+$\!\!\!\!\geq\!\!\!\!$ $(1-\delta_{UM})\|\mathcal{B}_{1}-\mathcal{B}_{2}\|_{F}^{2},$
 which guarantees the possibility of exact recovery of channel matrices in the channel tensor.
-Next, we present a formal analysis of the upper bound for ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ with ^$\widehat{\mathcal{B}}\in\mathbb{B}_{N,\mathcal{S}}$.
+Next, we present a formal analysis of the upper bound for $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ with $\widehat{\mathcal{B}}\in\mathbb{B}_{N,\mathcal{S}}$, obtained by minimizing (7), independent of the specific optimization algorithm.
 Here, we provide a concise summary of the main result.
 Theorem 2.
-(Upper bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}$ in (8) where $\bm{H}$ and $\bm{G}$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(LPN+UMN)}$ for a positive constant $c_{1}$, we have
-^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (12)
+(Upper bound of $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}$ in (8) where $\bm{H}$ and $\bm{G}$ are full rank, when $\mathcal{X}$ satisfies the $UM$-RIP and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(LPN+UMN)}$ for a positive constant $c_{1}$, the solution in (7) satisfies
+$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (13)
 where $\widehat{\mathcal{B}}$ is the solution to (7).
-The detailed analysis is provided in Appendix C.
+The detailed analysis is provided in the Appendix D.
 Theorem 2 guarantees a stable recovery of the ground truth $\mathcal{B}^{\star}$ when the number of time slots $T$ is linearly proportionate to $LPN+UMN$.
-Combining the condition $T\geq\Omega(UM/\delta_{UM}^{2})$ in (9) with ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
-$T\geq\max\bigg{\{}\Omega\bigg{(}\frac{UM}{\delta_{UM}^{2}}\bigg{)},\Omega\bigg{(}\frac{(1+\delta_{UM})(LPN+UMN)\gamma^{2}}{(1-\delta_{UM})^{2}\epsilon^{2}}\bigg{)}\bigg{\}}$
+For subgaussian measurements, combining the condition $T=\Omega(UM/\delta_{UM}^{2})$ in (10) with $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
+$T=\max\bigg\{\Omega\bigg(\frac{UM}{\delta_{UM}^{2}}\bigg),\Omega\bigg(\frac{(1+\delta_{UM})(LPN+UMN)\gamma^{2}}{(1-\delta_{UM})^{2}\epsilon^{2}}\bigg)\bigg\}$
 to ensure a small recovery error.
-Next we consider two special cases: $(i)$ Theorem 2 assumes that the channel matrices $\bm{H}$ and $\bm{G}$ are full rank, but in some scenarios, such as millimeter-wave MIMO systems, the presence of a large number of transmit/receive antennas combined with scattering-poor propagation may lead to low-rank channel matrices $\bm{H}$ and $\bm{G}$ [63, 15].
+On the other hand, when using the normalized DFT matrix, to guarantee $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq\epsilon$, the number of time slots $T$ should satisfy
+$T\geq\max\bigg\{UM,c\frac{(LPN+UMN)\gamma^{2}}{\epsilon^{2}}\bigg\},$
+where $c$ is a positive constant.
+We note that a commonly considered IRS-assisted MIMO scenario corresponds to $P=1$, i.e., a single base station, under which all theoretical results in this paper remain valid without any modification of the underlying assumptions.
+In addition, we emphasize that our result does not require any prior knowledge of the channel matrices $\bm{G}$ and $\bm{H}$ or the IRS configuration $\mathcal{S}$; it only relies on their dimensions.
+This generality naturally includes special scenarios such as channel hardening [63], which arises when $N$ is sufficiently large.
+Next we consider three special cases: $(i)$ Theorem 2 assumes that the channel matrices $\bm{H}$ and $\bm{G}$ are full rank, but in some scenarios, such as millimeter-wave MIMO systems, the presence of a large number of transmit/receive antennas combined with scattering-poor propagation may lead to low-rank channel matrices $\bm{H}$ and $\bm{G}$ [64, 15].
 When the signal travels between the BS and IRS across $r_{1}$ clusters, and between the IRS and the UT across $r_{2}$ clusters, it is reasonable to assume that $\bm{H}$ is a rank-$r_{1}$ matrix and $\bm{G}$ is a rank-$r_{2}$ matrix.
-By leveraging the covering number of low-rank matrices as introduced in [64, Lemma 3.1] and applying it to (40) in Appendix C, we can readily extend (12) to
-^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$
-$\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})((LP+N)r_{1}+(UM+N)r_{2})}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (13)
-where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N}\ \text{with}\ \text{rank}(\bm{H})$
-$=r_{1},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{with}\ \text{rank}(\bm{G})=r_{2},\text{are unknown and}$
+By leveraging the covering number of low-rank matrices as introduced in [65, Lemma 3.1] and applying it to (12) in Appendix D, we can readily extend (13) to
+$\hskip-34.14322pt\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ (14)
+$\hskip-34.14322pt=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})((LP+N)r_{1}+(UM+N)r_{2})}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
+where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N}\ \text{with}\ \text{rank}(\bm{H})$
+$r_{1},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{with}\ \text{rank}(\bm{G})=r_{2},\text{are unknown and}$
 $\mathcal{S}\in\mathbb{C}^{N\times K\times N}\ \text{is known}\}$.
-It is important to note that a comparison between (12) and (II) reveals that the improvement in recovery error due to low-rankness occurs only when $r_{1}\leq\frac{LPN}{LP+N}$ and $r_{2}\leq\frac{UMN}{UM+N}$.
-$(ii)$ It is crucial to emphasize that the elements of the IRS in outdoor scenarios, being exposed to weather and atmospheric conditions, may encounter unknown blockages and time-dependent fluctuations in their phase and amplitude responses [51].
+It is important to note that a comparison between (13) and (14) reveals that the improvement in recovery error due to low-rankness occurs only when $r_{1}\leq\frac{LPN}{LP+N}$ and $r_{2}\leq\frac{UMN}{UM+N}$.
+$(ii)$ In addition to low-rank structures, channel matrices in mmWave MIMO systems also exhibit sparsity in the angular domain due to the limited number of scattering paths [41, 32].
+In this case, $\bm{H}$ and $\bm{G}$ can be decomposed as $\bm{H}=\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2}$ and $\bm{G}=\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}$, where $\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}}$, $\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N}$, $\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}$, and $\bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}$ are known matrices defined in [41], and $\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}$ and $\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}$ are respectively unknown diagonal matrices containing only $s_{1}$ and $s_{2}$ nonzero entries corresponding to the path gains.
+When $s_{1}\leq\frac{F_{1}}{2}$ and $s_{2}\leq\frac{F_{2}}{2}$, the sparse channel structure admits an upper bound on the recovery error, given by:
+$\hskip-34.14322pt\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$
+$\hskip-34.14322pt=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(s_{1}\log(\frac{F_{1}}{s_{1}})+s_{2}\log(\frac{F_{2}}{s_{2}}))}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (15)
+where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}},\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}},\|\bm{\Sigma}_{1}\|_{0}=s_{1},\|\bm{\Sigma}_{2}\|_{0}=s_{2}\ \text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}},\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N},\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}\ \text{and}\ \bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}\ \text{are known}\}$.
+The detailed proof has been provided in the Appendix E.
+As shown in (II-C), when the logarithmic factor is ignored, the recovery error scales only with the sparsity levels $s_{1}$ and $s_{2}$, which explains the practical effectiveness of exploiting sparsity in mmWave MIMO systems.
+$(iii)$ It is crucial to emphasize that the elements of the IRS in outdoor scenarios, being exposed to weather and atmospheric conditions, may encounter unknown blockages and time-dependent fluctuations in their phase and amplitude responses [42].
 Consequently, the IRS tensor $\mathcal{S}$ deviates from its intended structure, and the assumption of perfect knowledge of all phase shifts at the receiver may not be valid in such cases.
-To analyze the scenario where $\mathcal{S}$ is unknown, by introducing the covering number for $\mathcal{S}$ to the proof in Appendix C , we can obtain
-^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\big{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+KN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\big{)},$
-where ^,$\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{are unknown}\}$.
+To analyze the scenario where $\mathcal{S}$ is unknown, by introducing the covering number for $\mathcal{S}$ to the proof in Appendix D, we can obtain
+$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\big(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+KN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\big),$
+where $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{G}\in\mathbb{C}^{N\times UM}\ \text{are unknown}\}$.
 Apart from the additional term $KN$, this upper bound closely aligns with (13).
@@
 Our second goal is to establish a minimax lower bound using Gaussian ensemble design.
-This lower bound proves to be a valuable tool for gaining insights into the parameters that influence the achievable error of a given problem.
-Furthermore, it reveals that the upper bound in (12) is tight and optimal.
-Following the derivation in Appendix D, we arrive at:
+This bound serves as a fundamental tool for understanding how the structural properties of a given model govern its achievable recovery error.
+Notably, this lower bound depends solely on the structure of the model and is independent of any specific loss function or estimator.
+Following the derivation in Appendix F, we arrive at:
 Theorem 3.
-(Minimax lower bound of ^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) In the context of the TT-based ToT regression model in (6), we assume a channel tensor $\mathcal{B}^{\star}$ from (8), where $\bm{H}$ and $\bm{G}$ are full rank and $N\geq C^{\prime}$ with $C^{\prime}$ being a universal constant.
-Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
-$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg{)}.$ (14)
-Note that the constraint $N\geq C^{\prime}$ is not strictly necessary, primarily simplifying the theoretical analysis in Appendix D.
+(Minimax lower bound of $\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$) In the context of the TT-based ToT regression model in (6), we assume a channel tensor $\mathcal{B}^{\star}$ from (8), where $\bm{H}$ and $\bm{G}$ are full rank and $N\geq C^{\prime}$ with $C^{\prime}$ being a universal constant.
+Assume that the noise $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, and that the pilot matrix $\bm{X}$ is chosen as either (i) an i.i.d.
+$\mathcal{C}\mathcal{N}(0,1)$ matrix or (ii) a normalized DFT matrix $\frac{1}{\sqrt{T}}\bm{X}$.
+Under these two choices, we obtain
+$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg).$ (16)
+Note that the constraint $N\geq C^{\prime}$ is not strictly necessary, primarily simplifying the theoretical analysis.
+By ignoring the constant term associated with $\delta_{UM}$ in (13), we conclude that the upper bound–independent of any specific optimization algorithm–on the recovery error of the constrained least squares minimization is tight and optimal.
+TABLE I: Comparison of Theoretical Guarantees: Ours vs. [31].
+Condition for $T$ Upper Bound on Error Lower Bound on Error
+[31] $K\cdot\min\{T,LP\}\geq N,T\geq UM$ (necessary) N/A $\Omega\big(\sqrt{\frac{LPNUM}{T}}\gamma\big)$
+Ours $T=\Omega\big(\frac{UM}{\delta_{UM}^{2}}\big)$ or $T\geq UM$ (sufficient) $O\big(\sqrt{\frac{(1+\delta_{UM})(LPN+UMN)}{(1-\delta_{UM})^{2}T}}\gamma\big)$ $\Omega\big(\sqrt{\frac{LPN+UMN}{T}}\gamma\big)$
+II-D Comparison with Prior Work [31]
+Similar to our theoretical result, [31] provides a detailed analysis of the number of time slots required to recover the channel matrices from (2), establishing that $K\cdot\min\{T,LP\}\geq N$ and $T\geq UM$ are necessary conditions.
+However, their analysis primarily concerns a specific bilinear alternating least squares (BALS) algorithm that minimizes a constrained least squares objective.
+In particular, the condition $K\cdot\min\{T,LP\}\geq N$ ensures the uniqueness of the solution in the CP decomposition, and $T\geq UM$ guarantees that $\bm{X}$ is left-invertible.
+In contrast, our results in (10) and (11), offer sufficient conditions for exact recovery under constrained least squares minimization in (7), independent of any specific algorithm.
+Moreover, [31] derives a specific lower bound for the CRB, which has been widely used as a benchmark for performance evaluation in IRS-assisted MIMO channel estimation [31, 34, 66, 67, 68, 69].
+Specifically, by defining $\bm{S}=\begin{bmatrix}\bm{s}(1)\cdots\bm{s}(K)\end{bmatrix}^{\top}\in\mathbb{C}^{K\times N}$, assuming $\frac{1}{T}\bm{X}\bm{X}^{\rm H}={\bf I}_{UM}$, and letting each entry of $\mathcal{W}$ follow the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, the following bound holds:
+$\operatorname{\mathbb{E}}\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}$ $\!\!\!\!=\!\!\!\!$ $\operatorname{\mathbb{E}}\|(\widehat{\bm{G}}^{\top}*\widehat{\bm{H}})\bm{S}^{\top}-(\bm{G}^{\top}*\bm{H})\bm{S}^{\top}\|_{F}$ (17)
+$\!\!\!\!=\!\!\!\!$ $\Omega\big(\sqrt{\frac{LPNUM}{T}}\gamma\big),$
+where the inequality follows from [31, eq.
+(70)].
+We note that (2) is inherently nonconvex, involving a total of $LPN+UMN$ degrees of freedom in the unknown channel matrices.
+In general, the recovery error is expected to scale with the number of degrees of freedom [70, 71, 72, 51, 50, 40].
+However, the bound in (17) scales only with $LPNUM$, as it is derived from a relaxed formulation that transforms the original nonconvex model in (2) into a convex one by treating $\bm{G}^{\top}*\bm{H}\in\mathbb{C}^{LPUM\times N}$ as a single unknown variable.As a result, this bound does not accurately reflect the complexity of the original nonconvex problem and is thus not optimal.
+In contrast, our work provides the first optimal recovery error bound for channel estimation in IRS-assisted MIMO systems that accounts for the full nonconvex structure of the problem.
+Recall that the minimax lower bound is given by
+$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}^{\star}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{T}}\gamma\bigg),$ (18)
+which scales with the total degrees of freedom, $LPN+UMN$, in the unknown channel matrices. A detailed comparison is summarized in Table I.
 III Generalization to Multi-Hop IRS-Assisted MIMO System
@@
 Furthermore, in the context of Terahertz (0.1-10 THz) band communication, considered a promising technology for achieving ultra-high speed and low-latency communications, deploying multiple passive IRSs between BSs and UTs proves effective in overcoming inherent propagation attenuations.
-Tensor-based model formulation
-In this section, we explore the $D$-IRS scheme connecting UTs and BSs, assuming unavailability of channels from UT to BS, from UT to the $i$-th IRS ($1<i\leq D$), from the $j$-th IRS ($1\leq j<D$) to BS, and between non-adjacent IRSs due to blockage or excessive path loss.
+III-A Tensor-based model formulation
+In this section, we explore the $D$-IRS scheme connecting UTs and BSs, assuming unavailability of channels from UTs to BSs, from UTs to the $i$-th IRS ($1<i\leq D$), from the $j$-th IRS ($1\leq j<D$) to BSs, and between non-adjacent IRSs due to blockage or excessive path loss.
 In detail, we denote the UT-$1$-th IRS channel as $\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM}$, the $D$-th IRS-BS channel as $\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}}$, the channel from the $d$-th IRS to the $(d+1)$-th IRS as $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ for $d\in[D-1]$, and represent the $d$-th IRS as $\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}}$ for $d\in[D]$.
 Here, $\mathcal{S}_{d}(:,k,:)=\text{diag}(\bm{s}_{d}(k))\in\mathbb{C}^{N_{d}\times N_{d}}$, and the design of $\bm{s}_{d}(k)$ is same with (2).
+Here, we emphasize that, to simplify our theoretical analysis, the IRSs are arranged in a cascaded structure where only the first IRS interacts directly with the UEs, and only the last IRS connects to the BSs.
+All intermediate IRSs function purely as passive relays and do not directly interface with any BS or UE.
+A visual illustration can be found in [77, Figure 1], where, upon excluding the direct path, the remaining configuration accurately represents our model.
 We define each element of the channel tensor $\mathcal{B}^{\star}_{D}\in\mathbb{C}^{LP\times K\times UM}$ as follows:
-[…]
+$\mathcal{B}^{\star}_{D}(p,k,j)$ $\!\!\!\!=\!\!\!\!$ $\bm{B}_{D}(p,:)\Pi_{d=1}^{D-1}\mathcal{S}_{d+1}(:,k,:)\bm{B}_{d}$ (19)
 $\cdot\mathcal{S}_{1}(:,k,:)\bm{B}_{0}(:,j).$
@@
 $\widetilde{\mathcal{Y}}(p,k,t)\!=\!\sum_{j=1}^{UM}\mathcal{B}^{\star}_{D}(p,k,j)\bm{X}(j,t)\!+\!\mathcal{W}(p,k,t),$ (20)
-where, in contrast to the multi-hop scheme [69], we assume the absence of a direct channel from the BS to the UT to streamline the analysis.
-Note that each element $\mathcal{B}^{\star}_{D}(p,k,j)$ can still be conceptualized as an element of an order-$2D+1$ TT format tensor due to its multiplication form resembling (29).
+where we differ from the multi-hop scheme in [77] by assuming the absence of a direct BS-UT channel.
+In addition, we neglect both direct and reflected interference across distinct BS-UT links, effectively treating all paths as isolated.
+Furthermore, signal propagation and interference between different IRSs are neglected.
+These assumptions effectively decouple the communication paths, allowing us to concentrate on the highly nonconvex structure of the model in (19).
+To the best of our knowledge, a rigorous theoretical characterization of the recovery error for this model remains absent in the literature.
+Note that each element $\mathcal{B}^{\star}_{D}(p,k,j)$ can still be conceptualized as an element of an order-$2D+1$ TT format tensor due to its multiplication form resembling (1) in the Appendix A.
 Specifically, to establish a connection between the channel tensor $\mathcal{B}^{\star}_{D}$ and the TT format tensor, we can reinterpret $\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}}$ as an order-$3$ tensor $\widetilde{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times 1\times N_{d}}$.
-Accordingly, we define ^$\widehat{\mathcal{B}}_{D}^{\star}$ as the result of a sequence of tensor contractions  [^4] Here, the tensor contraction operation $\mathcal{A}_{1}\times_{d}^{1}\mathcal{A}_{2}$ for $\mathcal{A}_{1}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ and $\mathcal{A}_{2}\in\mathbb{C}^{N_{1}\times M_{2}\times M_{3}}$ results in a new tensor $\mathcal{A}_{3}$ of size $N_{1}\times\cdots\times N_{d-1}\times N_{d+1}\times\cdots\times N_{D}\times M_{2}\times M_{3}$, with the $(n_{1},\dots,n_{d-1},n_{d+1},\dots,n_{D},m_{2},m_{3})$-th element being $\sum_{n_{d}}\mathcal{A}_{1}(n_{1},\dots,n_{d})\mathcal{A}_{2}(n_{d},m_{2},m_{3})$.: $\bm{B}_{D}\times_{2}^{1}\mathcal{S}_{D}\times_{3}^{1}\widetilde{\bm{B}}_{D-1}\times_{4}^{1}\cdots\times_{2D-1}^{1}\widetilde{\bm{B}}_{1}\times_{2D}^{1}\mathcal{S}_{1}\times_{2D+1}^{1}\bm{B}_{0}$.
-As a result, ^$\widehat{\mathcal{B}}_{D}^{\star}$ can be represented as an order-$2D+1$ TT format tensor with dimensions $LP\times K\times 1\times K\times\cdots\times 1\times K\times UM$ with TT ranks  [^5] The definition of TT ranks is provided in Appendix A.
+Accordingly, we define $\widehat{\mathcal{B}}_{D}^{\star}$ as the result of a sequence of tensor contractions [^3] Here, the tensor contraction operation $\mathcal{A}_{1}\times_{d}^{1}\mathcal{A}_{2}$ for $\mathcal{A}_{1}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ and $\mathcal{A}_{2}\in\mathbb{C}^{N_{1}\times M_{2}\times M_{3}}$ results in a new tensor $\mathcal{A}_{3}$ of size $N_{1}\times\cdots\times N_{d-1}\times N_{d+1}\times\cdots\times N_{D}\times M_{2}\times M_{3}$, with the $(n_{1},\dots,n_{d-1},n_{d+1},\dots,n_{D},m_{2},m_{3})$-th element being $\sum_{n_{d}}\mathcal{A}_{1}(n_{1},\dots,n_{d})\mathcal{A}_{2}(n_{d},m_{2},m_{3})$.: $\bm{B}_{D}\times_{2}^{1}\mathcal{S}_{D}\times_{3}^{1}\widetilde{\bm{B}}_{D-1}\times_{4}^{1}\cdots\times_{2D-1}^{1}\widetilde{\bm{B}}_{1}\times_{2D}^{1}\mathcal{S}_{1}\times_{2D+1}^{1}\bm{B}_{0}$.
+As a result, $\widehat{\mathcal{B}}_{D}^{\star}$ can be represented as an order-$2D+1$ TT format tensor with dimensions $LP\times K\times 1\times K\times\cdots\times 1\times K\times UM$ with TT ranks [^4] The definition of TT ranks is provided in the Appendix A.
 $(N_{D},N_{D},N_{D-1},N_{D-1},\dots,N_{1},N_{1})$.
 Further insights can be gained, revealing that
-$\mathcal{B}^{\star}_{D}(p,k,j)$
-={^$=\begin{cases}\widehat{\mathcal{B}}_{D}^{\star}(p,k_{1},1,k_{2},1,\dots,k_{D},j),&k_{1}=\dots=k_{D}=k,\\ 0,&\text{otherwise},\end{cases}$ (17)
-where it becomes evident that $\mathcal{B}^{\star}_{D}$ emerges as a sampled outcome of ^$\widehat{\mathcal{B}}_{D}^{\star}$.
+$\hskip-22.76228pt\mathcal{B}^{\star}_{D}(p,k,j)$
+$\hskip-22.76228pt=\begin{cases}\widehat{\mathcal{B}}_{D}^{\star}(p,k_{1},1,k_{2},1,\dots,k_{D},j),&k_{1}=\dots=k_{D}=k,\\ 0,&\text{otherwise},\end{cases}$ (21)
+where it becomes evident that $\mathcal{B}^{\star}_{D}$ emerges as a sampled outcome of $\widehat{\mathcal{B}}_{D}^{\star}$.
 To streamline the notation, we introduce the notation $\mathcal{B}^{\star}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}$.
 Consequently, we define the TT-based ToT regression model as following:
-[…]
-[…]
+$\widetilde{\mathcal{Y}}$ $\!\!\!\!=\!\!\!\!$ $\mathcal{X}(\mathcal{B}^{\star}_{D})+\mathcal{W}=\mathcal{B}^{\star}_{D}\times_{3}^{1}\bm{X}+\mathcal{W}$ (22)
+$\!\!\!\!=\!\!\!\!$ $[\bm{B}_{D},\mathcal{S}_{D},\bm{B}_{D-1},\dots,\mathcal{S}_{1},\bm{B}_{0}]\times_{3}^{1}\bm{X}+\mathcal{W}.$
+Here, we emphasize that the TT-based ToT regression model is equivalent to (20), and serves to simplify the subsequent error analysis.
 Now, we consider the following constrained least squares objective:
-^$\widehat{\mathcal{B}}_{D}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})-\widetilde{\mathcal{Y}}\|_{F}^{2},$ (19)
+$\widehat{\mathcal{B}}_{D}=\operatorname*{\text{arg~min}}_{\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})-\widetilde{\mathcal{Y}}\|_{F}^{2},$ (23)
 where we define the set of $\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$ as follows:
-[…]
+$\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$ $\!\!\!\!\!\!=\!\!\!\!\!\!$ $\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:$ (24)
 $\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},$
@@
 in which $N=\max_{d}N_{d}$. Note that each element of $\mathcal{B}_{D}$ follows (19). In the subsequent section, we present a comprehensive error analysis of the channel estimation process.
-Error analysis
+III-B Optimal error analysis
 Before presenting the error analysis, we first ensure that the measurements satisfy the RIP condition.
-By following a similar derivation as provided in Appendix B, the RIP condition for $\mathcal{X}(\mathcal{B}_{D})$ can be directly obtained.
+By following a similar derivation as provided in Appendix C, the RIP condition for $\mathcal{X}(\mathcal{B}_{D})$ can be directly obtained.
 Theorem 4.
-Suppose each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable.
-Let $\delta_{UM}\in(0,1)$ be a positive constant.
-Then, for any channel tensor $\mathcal{B}^{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$, when the number of time slots satisfies
+A linear operator $\mathcal{X}:\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ is said to satisfy the $UM$-RIP with a constant $\delta_{UM}$ if
+$(1-\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2}\!\leq\!\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})\|_{F}^{2}\!\leq\!(1+\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2},$ (25)
+holds for any channel tensor $\mathcal{B}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}$. We now consider two cases:
+  • Suppose each element of the pilot signal matrix $\bm{X}$ is an i.i.d.
+complex-valued subgaussian random variable.
+For a positive constant $\delta_{UM}\in(0,1)$, when the number of time slots satisfies
     $T\geq C\cdot\frac{UM}{\delta_{UM}^{2}},$ (26)
-with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP:
-$(1-\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2}\!\leq\!\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})\|_{F}^{2}\!\leq\!(1+\delta_{UM})\|\mathcal{B}_{D}\|_{F}^{2},$ (22)
-where $c$ and $C$ is a universal constant.
+    with probability $1-e^{-cT}$, $\mathcal{X}$ satisfies the $UM$-RIP with positive constants $c$ and $C$.
+  • Suppose $\frac{1}{\sqrt{T}}\bm{X}$ is the normalized DFT matrix, when
+    $T\geq UM$ (27)
+    is satisfied, we have $\frac{1}{T}\|\mathcal{X}(\mathcal{B}_{D})\|_{F}^{2}\!=\!\|\mathcal{B}_{D}\|_{F}^{2}$ with $\delta_{UM}=0$.
 Similar to Theorem 1, Theorem 4 ensures that the number of time slots is solely contingent upon the total number of antennas (UM) in the UTs.
 This is attributed to the fact that the pilot signal matrix $\bm{X}$ is exclusively associated with the UT-$1$-th IRS channel $\bm{B}_{0}$.
-Then we can further analyze the recovery error ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$.
+Then we can further analyze the recovery error $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$.
 Theorem 5.
-(Upper bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}_{D}$ in (20) in which $\bm{B}_{d},d=0,\dots,D$ are full rank, when each element of the pilot signal matrix $\bm{X}$ is a complex-valued subgaussian random variable and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}$ for a positive constant $c_{1}$, we have
-^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg{)},$ (23)
-where $N_{0}=UM$, $N_{D+1}=LP$ and  ^$\widehat{\mathcal{B}}_{D}$ is the solution to (19).
-The proof has been provided in Appendix E.
-Theorem 5 ensures an optimal recovery error bound, given that the degree of freedom for the unknown variables in (19) is $O(LPN_{D}+UMN_{1}+\sum_{i=1}^{D-1}N_{d}N_{d+1})$.
+(Upper bound of $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) Given a channel tensor $\mathcal{B}^{\star}_{D}$ in (24) in which $\bm{B}_{d},d=0,\dots,D$ are full rank, when $\mathcal{X}$ satisfies the $UM$-RIP and each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, with probability $1-2e^{-c_{1}(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}$ for a positive constant $c_{1}$, the solution in (23) satisfies
+$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$ (28)
+where $N_{0}=UM$, $N_{D+1}=LP$ and  $\widehat{\mathcal{B}}_{D}$ is the solution to (23).
+The proof has been provided in Appendix G.
+Theorem 5 provides an upper bound on the recovery error derived from a constrained least squares minimization, which scales with the total degrees of freedom of the unknown variables, i.e., $LPN_{D}+UMN_{1}+\sum_{i=1}^{D-1}N_{d}N_{d+1}$.
 However, it is noteworthy that while increasing the number of IRSs can expand service coverage, the recovery error of channel estimation may also increase when the number of time slots $T$ is limited.
 Therefore, in practical implementation, a balance should be struck between service coverage and recovery error.
-Similar to the analysis of the minimax lower bound using the Gaussian ensemble design in Theorem 3, this result can also be extended to multi-hop IRS-assisted MIMO systems.
+A discussion of special cases analogous to the single-hop IRS-assisted MIMO system is presented in Appendix H.
+Building on the upper bound derived from constrained least squares minimization, we now establish a minimax lower bound to assess its optimality.
 Theorem 6.
-(Minimax lower bound of ^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) In the context of the TT-based ToT regression model in (18), we assume a channel tensor $\mathcal{B}^{\star}_{D}$ from (20), where $\bm{B}_{d},d=0,\dots,D$ are full rank and $\min N_{d}\geq C^{\prime}$ with $C^{\prime}$ being a universal constant.
-Assuming each element of $\bm{X}$ and $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, respectively, we obtain:
-$\inf_{\widehat{\mathcal{B}}_{D}}\sup_{\mathcal{B}^{\star}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{\sum_{d=0}^{D}N_{d}N_{d+1}}{T}}\gamma\bigg{)}.$ (24)
-Finally, akin to the analysis of a single-hop IRS-assisted MIMO system, it is imperative to address the following two special cases.
-$(i)$ First, when $\bm{B}_{d},d=0,\dots D$ are rank-$r_{d}$ matrices, by incorporating the covering number of low-rank matrices into (60) in Appendix E, we can derive
-^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
-[…]
-where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]\ \text{are unknown with }\text{rank}(\bm{B}_{d})=r_{d},d=0,\dots D,\ \text{and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$.
-Leveraging low-rank structures offers a potential avenue for improving error performance, as the recovery error may be reduced below (23) when the ranks $r_{d},d=0,\dots,D$ are small.
-$(ii)$ Second, when $\mathcal{S}_{i},i\in[D]$, are unknown, introducing the covering number of $\{\mathcal{S}_{i}\}_{i=1}^{D}$ into the proof in Appendix E, similar to (23), we can arrive at
-^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
-[…]
-where ^$\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{N_{D+1}\times K\times N_{0}}:\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d=0,\dots,D,\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are unknown}\}$.
-In this analysis, we assume that all IRSs are equipped with active elements.
-However, in practice, when adjustments are required for only a subset of IRSs, denoted as $\mathcal{S}_{i}$ for $i\in\Omega$, where $\Omega\subseteq\{1,\dots,D\}$, the recovery error can be further reduced.
-Specifically, the term $\sum_{d=1}^{D}KN_{d}$ is replaced by $\sum_{i\in\Omega}KN_{i}$.
-Discussion
+(Minimax lower bound of $\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$) In the context of the TT-based ToT regression model in (22), we assume a channel tensor $\mathcal{B}^{\star}_{D}$ from (24), where $\bm{B}_{d},d=0,\dots,D$ are full rank and $\min N_{d}\geq C^{\prime}$ with $C^{\prime}$ being a universal constant.
+Assume that the noise $\mathcal{W}$ in (6) follows $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, and that the pilot matrix $\bm{X}$ is chosen as either (i) an i.i.d.
+$\mathcal{C}\mathcal{N}(0,1)$ matrix or (ii) a normalized DFT matrix $\frac{1}{\sqrt{T}}\bm{X}$.
+Under these two choices, we obtain
+$\inf_{\widehat{\mathcal{B}}_{D}}\sup_{\mathcal{B}^{\star}_{D}\in\mathbb{B}_{N,\{\mathcal{S}_{i}\}}^{D}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}=\Omega\bigg(\sqrt{\frac{\sum_{d=0}^{D}N_{d}N_{d+1}}{T}}\gamma\bigg).$ (29)
+Our minimax lower bound for the TT-based ToT regression model in (22) demonstrates that the recovery error upper bound is tight and optimal up to the log term.
+Figure 2: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system (a) for different NN with $UM=LP=K=10$, $T=20$ and $\gamma^{2}=10^{-6}$, (b) for different $T$ with $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (c) for different $\gamma^{2}$ with $UM=LP=K=N=10$ and $T=20$.
+III-C Discussion
 Beyond the IRS-assisted MIMO system, the tensor-based ToT model can also be effectively applied to a variety of typical communication systems, including direct-sequence code-division multiple-access systems [80], space-time frequency MIMO systems [81], and cooperative/relay systems [82, 83].
@@
 These results offer a theoretical guarantee for robustness and potential accuracy in estimating channel characteristics.
-IV Nonconvex Optimization Algorithms
-In this section, our focus will be on designing algorithms for estimating the channel matrices within the framework of the TT-based ToT regression model.
-Initially, we will concentrate on the single-hop IRS-assisted MIMO system and subsequently extend our approach to the multi-hop IRS-assisted MIMO system.
-Note that, given that the following algorithms can be readily extended to handle the unknown $\mathcal{S}$ scenario, our attention is directed towards the known $\mathcal{S}$ case at the receiver.
-Single-hop IRS-assisted MIMO system
-$(i)$ One common assumption is that the pilot signal matrix $\bm{X}$ is a truncated discrete Fourier transform (DFT) matrix or a semi-unitary matrix, as discussed in previous works [13, 31].
-However, we can extend this assumption by considering the existence of the pseudoinverse for $\bm{X}$.
-Expanding upon the equivalent form of the TT format $\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]$ [76], denoted as $\mathcal{B}^{\langle 1\rangle}=\bm{H}R(\mathcal{S})(\bm{G}\otimes{\bf I}_{K})$ and $\mathcal{B}^{\langle 2\rangle}=({\bf I}_{K}\otimes\bm{H})L(\mathcal{S})\bm{G}$ (The expressions for $L(\mathcal{S})$ and $R(\mathcal{S})$ are provided in (31) and (32) of Appendix A.)  [^6] Let $\mathcal{X}^{\langle i\rangle}\in\mathbb{C}^{(d_{1}\cdots d_{i})\times(d_{i+1}\cdots d_{N})}$ denote the $i$-th unfolding matrix of the tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times\dots\times d_{N}}$.
-The $(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})$-th element of $\mathcal{X}^{\langle i\rangle}$ is given by $\mathcal{X}^{\langle i\rangle}(s_{1}\cdots s_{i},s_{i+1}\cdots s_{N})=\mathcal{X}(s_{1},\dots,s_{N})$.
-Here, $s_{1}\cdots s_{i}$ and $s_{i+1}\cdots s_{N}$ represent the row and column indices, respectively, where the row index is calculated as $s_{1}+d_{1}(s_{2}-1)+\cdots+d_{1}\cdots d_{i-1}(s_{i}-1)$ and the column index as $s_{i+1}+d_{i+1}(s_{i+2}-1)+\cdots+d_{i+1}\cdots d_{N-1}(s_{N}-1)$., we can formulate the following loss function:
-$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}}\|\widehat{\bm{H}}R(\mathcal{S})(\widehat{\bm{G}}\otimes{\bf I}_{K})-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}\|_{F}^{2},$ (25)
-$\min_{\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}\|({\bf I}_{K}\otimes\widehat{\bm{H}})L(\mathcal{S})\widehat{\bm{G}}-(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}\|_{F}^{2}.$ (26)
-Specifically, we apply the following alternating least squares (ALS) algorithm:
-$\bm{H}^{(l+1)}=(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 1\rangle}(R(\mathcal{S})(\bm{G}^{(l)}\otimes{\bf I}_{K}))^{\dagger},$
-$\bm{G}^{(l+1)}=(({\bf I}_{K}\otimes\bm{H}^{(l+1)})L(\mathcal{S}))^{\dagger}(\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger})^{\langle 2\rangle}.$
-To guarantee the existence of all pseudoinverse matrices, it is imperative that the conditions $N\leq\min\{UMK,LPK\}$ and $T\geq UM$ are met but we note that $N\leq\min\{UMK,LPK\}$ is a sufficient condition, not a necessary one.
-This is because Equation 12 holds for all coefficients when $T\geq\Omega(UM)$.
-Additionally, when $\bm{H}$ and $\bm{G}$ exhibit low-rank characteristics, we can employ SVD to truncate small singular values, facilitating the use of low-rank projection in each iteration.
-$(ii)$ To address the numerical stability challenges of the pseudoinverse and to accommodate a broader class of pilot signal matrices $\bm{X}$, we can alternatively optimize the following loss function directly:
-$\min_{\widehat{\bm{H}}\in\mathbb{C}^{LP\times N}\atop\widehat{\bm{G}}\in\mathbb{C}^{N\times UM}}f(\widehat{\bm{H}},\widehat{\bm{G}})=\frac{1}{T}\|\mathcal{X}([\widehat{\bm{H}},\mathcal{S},\widehat{\bm{G}}])-\mathcal{Y}\|_{F}^{2}.$ (27)
-As for (27), we solve the above optimization problem by the following alternating gradient descent (AGD):
-$\bm{H}^{(l+1)}=\bm{H}^{(l)}-\mu\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)}),$
-$\bm{G}^{(l+1)}=\bm{G}^{(l)}-\mu\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)}),$
-where $\mu$ is a step size and we define the Wirtinger gradients $\nabla_{\bm{H}^{*}}f(\bm{H}^{(l)},\bm{G}^{(l)})$ and $\nabla_{\bm{G}^{*}}f(\bm{H}^{(l+1)},\bm{G}^{(l)})$ as follows:
-[…]
-×$\times_{2,3}^{2,3}(\mathcal{S}^{*}\times_{3}^{1}(\bm{G}^{(l)})^{*}),$
-[…]
-×$\times_{1,2}^{1,2}\big{(}(\mathcal{B}^{(l)}\times_{3}^{1}\bm{X}-\mathcal{Y})\times_{3}^{2}\bm{X}^{*}\big{)}.$
-In comparison to the ALS algorithm, the AGD does not require additional coefficient constraints, aside from the condition $T\geq\Omega(UM)$ for satisfying the RIP condition.
-Additionally, it avoids the numerical stability issues associated with the pseudoinverse.
-Figure 1: Recovery performance for channel matrices (a) for different $T$ with $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (b) for different $UM$ with $T=100$, $LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (c) for different $LP$ with $T=20$, $UM=K=N=10$ and $\gamma^{2}=10^{-6}$.
-Figure 2: Recovery performance for channel matrices (a) for different $N$ with $T=20$, $UM=LP=K=10$ and $\gamma^{2}=10^{-6}$, (b) for different $K$ with $T=20$, $UM=LP=N=10$ and $\gamma^{2}=10^{-6}$, $\mu=0.02$ for the $1$ IRS scenario and $\mu=0.1$ for the $2$ IRSs scenario, (c) for different $\gamma^{2}$ with $T=20$ and $UM=LP=K=N=10$.
-Multi-hop IRS-assisted MIMO system
-In a multi-hop IRS-assisted MIMO system, where channel matrices and IRSs are represented in (20), a closed-form solution for the ALS method is not feasible.
-Consequently, we primarily consider the AGD method and the following loss function:
-$\operatorname*{\text{arg~{}min}}_{\widehat{\bm{B}}_{D}\in\mathbb{C}^{LP\times N_{D}},\widehat{\bm{B}}_{0}\in\mathbb{C}^{N_{1}\times UM},\atop\widehat{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]}$ $g(\widehat{\bm{B}}_{D},\dots,\widehat{\bm{B}}_{0})$ (28)
-[…]
-We solve this nonconvex optimization problem by the following AGD algorithm for $d=0,\dots,D$:
-$\bm{B}_{d}^{(l+1)}=\bm{B}_{d}^{(l)}\!\!-\!\mu\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D}^{(l+1)},\dots,\bm{B}_{d+1}^{(l+1)},\bm{B}_{d}^{(l)},\dots,\bm{B}_{0}^{(l)}),$
-where $\mu$ is a step size and the Wirtinger gradient $\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\\ \dots,\bm{B}_{0})$ is defined as:
-$\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})=\frac{1}{T}\sum_{k=1}^{K}\mathcal{S}_{d+1}^{\rm H}(:,k,:)\bm{B}_{d+1}^{\rm H}\cdots$
-$\hskip 14.22636pt\mathcal{S}_{D}^{\rm H}(:,k,:)\bm{B}_{D}^{\rm H}\bigg{(}\mathcal{X}([\bm{B}_{D},\mathcal{S}_{D}(:,k,:),\dots,\bm{B}_{1},\mathcal{S}_{1}(:,k,:),$
-$\hskip 14.22636pt\bm{B}_{0}])-\widetilde{\mathcal{Y}}(:,k,:)\bigg{)}\bm{X}^{\rm H}\bm{B}_{0}^{\rm H}\cdots\bm{B}_{d-1}^{\rm H}\mathcal{S}_{d}^{\rm H}(:,k,:).$
-In line with the single-hop IRS-assisted MIMO system, AGD in this scenario similarly operates without coefficient constraints, except for the condition $T\geq\Omega(UM)$ for satisfying the RIP condition.
-We note that for low-rank channel estimation, iterative hard thresholding combined with truncated singular value decomposition [77] can be effectively integrated into the AGD method.
-V Experimental Results
-In this section, we conduct numerical experiments to evaluate the effectiveness of Theorem 2 and Theorem 5.
-We consider $1$ IRS and $2$ IRSs assisted MIMO systems, where each element of channel matrices $\bm{H}$, $\bm{G}$, $\bm{B}_{0}$, $\bm{B}_{1}$, $\bm{B}_{2}$, and pilot signal matrix $\bm{X}$ follows independent and identically complex standard normal distribution.
-Subsequently, we normalize the above channel matrices to have a unit Frobenius norm.
-To effectively estimate the channel matrices in the $1$ IRS assisted MIMO system, we apply the ALS, AGD, tensor train singular value decomposition (TT-SVD) [37], and tensor train cross-approximation (TT CA) via the greedy restricted cross-interpolation (GRCI) algorithm [78].
-As two representative methods for TT decomposition, TT-SVD and TT-CA are employed for the tensor factorization of $\mathcal{Y}\times_{3}^{1}\bm{X}^{\dagger}$.
-In the case of the $2$ IRSs model, we employ the AGD algorithm.
-We initialize ALS and AGD with random matrices, with elements generated from a complex standard normal distribution.
-To simplify parameter selection and notation in the $2$ IRSs system, we set $N_{1}=N_{2}=N$ and denote $\mathcal{B}_{2}=\mathcal{B}$, $\mathcal{B}_{2}^{\star}=\mathcal{B}^{\star}$.
-The number of iterations for the ALS, AGD ($1$ IRS), and AGD ($2$ IRSs) algorithms are set to $100$, $10^{5}$, and $2\times 10^{5}$, respectively.
-For the AGD algorithm, we set $\mu=0.2$ for the $1$ IRS scenario and $\mu=1$ for the $2$ IRSs scenario.
-For each experimental setting, we conduct 20 Monte Carlo trials and then take the average over the 20 trials to report the results.
-In the first experiment, we compare the recovery performance of different methods with various values of $T$, $UM$, and $LP$ in Figure 1.
-These plots illustrate that the recovery error tends to decrease with larger values of $T$, or smaller values of $UM$ and $LP$, consistent with our theoretical findings presented in Theorem 2 and Theorem 5.
-Additionally, we observe the following: $(i)$ when $T$ approaches $UM$, the recovery error could worsen since RIP conditions (9) and (21) are not satisfied; $(ii)$ AGD (1 IRS) and ALS exhibit similar recovery performance except for when $T=UM$; $(iii)$ the recovery error of AGD (2 IRSs) is larger than AGD (1 IRS) due to the introduction of an additional channel matrix, leading to more error as shown in Theorem 5; $(iv)$ compared to the ALS and AGD algorithms, TT-SVD and TT CA demonstrate poorer recovery performance.
-This is because TT-SVD and TT CA do not utilize the known information of $\mathcal{S}$ or $\mathcal{S}_{i},i=1,2$, resulting in a larger error.
-Furthermore, the recovery error of TT-CA is larger than that of TT-SVD, as TT-SVD utilizes the full information of the channel tensor, whereas TT-CA does not [79]; $(v)$ from Figure 1, it is evident that successful recovery is not dependent on $LP$, as even with $LP=100$, the recovery remains stable when $T=20$.
-In the second experiment, we examine the recovery performance of different methods across varying values of $N$ and $K$.
-Keeping a fixed number of time slots $T=20$, we observe stable recovery errors with increasing $N$ and $K$.
-However, it is noteworthy that as $K$ increases, the recovery performance generally improves, except for TT-SVD and TT CA.
-This is attributed to the fact that the coherence time $T_{s}=KT$ becomes larger as $K$ increases while $T$ remains fixed, thereby enhancing performance.
-However, factorization approaches like TT-SVD and TT CA struggle to effectively denoise as $K$ increases.
-This is due to the possibility that the estimated $\widehat{\mathcal{S}}$ or $\widehat{\mathcal{S}}_{1},\widehat{\mathcal{S}}_{2}$ may introduce more noise under these conditions.
-This highlights a limitation of tensor factorization in recovering channel matrices of IRS-based systems.
-In the third experiment, we compare the recovery performance of different methods across various noise variances $\gamma^{2}$.
-In Figure 2, we observe that, except for TT CA, the recovery error exhibits a linear relationship with $\gamma^{2}$, consistent with the analysis in Theorem 2 and Theorem 5.
-Ultimately, while the ALS achieves performance similar to the AGD, the AGD may offer greater stability, especially in time-varying systems where the pseudoinverse of the input matrix may not be well-defined.
-Figure 3: Recovery performance for channel matrices for different $D$ with $K=N_{0}=\cdots N_{D}=10$, $\gamma^{2}=10^{-6}$ and $\mu$ sequentially selected from the set $\{0.2,1,5,10,15\}$.
+IV Experimental Results
+In this section, we present numerical experiments to validate the effectiveness of the bounds in Theorems 2, 3, 5 and 6 for single-hop and multi-hop IRS-assisted MIMO systems [^5] Here, we replace $\delta_{UM}$ with $c\sqrt{\frac{UM}{T}}$, where $c$ is a positive constant, in accordance with (10) and (26).
+Moreover, since the performance metric in these experiments is the normalized recovery error $\frac{\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}^{2}}{\|\mathcal{B}^{\star}\|_{F}^{2}}$, we accordingly normalize the upper and lower bounds in Theorems 2, 3, 5 and 6 by dividing them by $\|\mathcal{B}^{\star}\|_{F}^{2}$..
+To simplify parameter selection, each entry of the channel matrices and the pilot signal matrix is independently drawn from the standard complex normal distribution.
+The channel matrices are then normalized to have unit Frobenius norm.
+Additionally, we set $N_{1}=\cdots=N_{D-1}=N$ and denote $\mathcal{B}_{D}=\mathcal{B}$ and $\mathcal{B}_{D}^{\star}=\mathcal{B}^{\star}$ for notational simplicity.
+To effectively estimate the channel matrices in the single-hop IRS-assisted MIMO system, we adopt the bilinear alternating least squares (BALS) algorithm [31], which is designed for constrained least squares minimization and has been widely used as a benchmark in channel estimation [32, 33, 84, 85, 86].
+In contrast to BALS, which admits closed-form updates in the single-hop case, the constrained least squares problem in the multi-hop IRS-assisted MIMO system ($D\geq 2$) lacks an analytical solution.
+To address this, we adopt the alternating gradient descent (AGD) algorithm with learning rate $\mu$, following the gradient-based strategy proposed in [77].
+The implementation details and computational complexities of both algorithms are provided in the Appendix B.
+We initialize BALS and AGD with random matrices, with elements generated from a complex standard normal distribution.
+The number of iterations is set to $100$ for BALS and $10000$ for AGD.
+For each experimental setting, we perform 20 Monte Carlo trials and report the average results across these trials.
+In the first experiment, we investigate the recovery performance of the BALS algorithm under varying values of $UM$ and $LP$, as depicted in Figure 1.
+The result illustrates that the recovery error consistently falls within the theoretical range defined by the upper and lower bounds in Theorems 2 and 3.
+In particular, the observed recovery error closely matches the upper bound associated with constrained least squares minimization, which scales proportionally with $LPN+UMN$, as characterized in Theorem 2.
+This scaling behavior reflects the intrinsic nonconvex structure of the IRS-assisted MIMO channel estimation problem, where $LPN$ and $UMN$ correspond to the degrees of freedom for the two channel matrices involved in the estimation problem.
+The empirical agreement with the theoretical bound further confirms the tightness and optimality of our recovery guarantees.
+Figure 1: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system for different UMUM and $LP$ with $K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$.
+In the second experiment, we systematically examine the recovery performance of the BALS algorithm across varying values of $N$, $T$ and $\gamma^{2}$.
+The empirical findings illustrated in Figure 2 demonstrate strong concordance with our theoretical guarantees.
+Notably, the recovery error exhibits an increasing trend with larger $N$ and $\gamma^{2}$, and a decreasing trend as the number of time slots $T$ grows, reflecting the fundamental trade-offs inherent in the estimation problem.
+Crucially, the observed recovery errors are rigorously confined within the analytically derived upper and lower bounds specified in Theorems 2 and 3, thereby substantiating the tightness and validity of our theoretical framework.
+Figure 3: Recovery performance of channel matrices in the single-hop IRS assisted MIMO system for different $\gamma_{1}^{2}$ with $UM=LP=K=N=10$, $\gamma^{2}=10^{-6}$ and $T=100$.
+In the preceding analysis, the BS-UT direct channels were omitted when evaluating the recovery performance in the single-hop IRS-assisted MIMO system.
+However, such direct links are often present in practical deployments.
+To account for this, the model in (2) can be reformulated as $\bm{Y}(k)=\bm{H}\text{diag}(\bm{s}(k))\bm{G}\bm{X}+\bm{A}\bm{X}+\bm{W}(k),k\in[K]$, where $\bm{A}\in\mathbb{C}^{LP\times UM}$ represents the BS-UT channel matrix.
+In the third experiment, we investigate the impact of the BS-UT channel $\bm{A}$ on the recovery performance of the BALS algorithm.
+Importantly, the recovery objective remains focused on estimating $\bm{H}$ and $\bm{G}$, rather than jointly estimating $\bm{H}$, $\bm{G}$, and $\bm{A}$.
+We further assume that each entry of $\bm{A}$ follows $\mathcal{C}\mathcal{N}(0,\gamma_{1}^{2})$.
+As shown in Figure 3, the recovery error increases as $\gamma_{1}^{2}$ becomes larger.
+This behavior arises because the term $\bm{A}\bm{X}$ effectively acts as an additional noise component when optimizing $\bm{H}$ and $\bm{G}$, similar in effect to $\bm{W}(k)$.
+Specifically, the $(i,j)$-th entry of $\bm{A}\bm{X}+\bm{W}(k)$ follows a complex Gaussian distribution $\mathcal{C}\mathcal{N}(0,\gamma_{1}^{2}\|\bm{X}(:,j)\|_{2}^{2}+\gamma^{2})$.
+As a result, the conclusions in Theorems 2 and 3 can be naturally extended by replacing $\gamma^{2}$ with $\max_{j}\gamma_{1}^{2}\|\bm{X}(:,j)\|_{2}^{2}+\gamma^{2}$.
+The empirical recovery errors remain tightly bounded by the theoretically derived upper and lower limits, thus confirming the validity of the theoretical analysis.
+In the fourth experiment, we evaluate the recovery performance of the AGD algorithm in multi-hop IRS-assisted MIMO systems with $D=2$, under varying values of $UM$, LPLP, NN, $T$ and $\gamma^{2}$.
+As shown in Figures 4, 4, 4, 5 and 5, the observed recovery errors exhibit trends that are consistent with the theoretical predictions in Theorems 5 and 6.
+Moreover, the empirical recovery errors remain tightly confined within the theoretically established upper and lower bounds, thereby providing additional evidence for the optimality of the derived recovery guarantees.
+Figure 4: Recovery performance of channel matrices in the multi-hop IRS assisted MIMO system (a) for different UMUM with $D=2$, $\mu=1$, $LP=K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$, (b) for different LPLP with $D=2$, $\mu=1$, $UM=K=N=10$, $T=100$ and $\gamma^{2}=10^{-6}$, (c) for different NN with $D=2$, $\mu=1$, $T=UM=LP=K=10$ and $\gamma^{2}=10^{-6}$.
+Figure 5: Recovery performance of channel matrices in the multi-hop IRS assisted MIMO system (a) for different $T$ with $D=2$, $\mu=1$, $UM=LP=K=N=10$ and $\gamma^{2}=10^{-6}$, (b) for different $\gamma^{2}$ with $D=2$, $\mu=1$, $UM=LP=K=N=10$ and $T=20$, (c) for different $D$ with $T=UM=LP=K=N=10$, $\gamma^{2}=10^{-6}$ and $\mu$ sequentially selected from the set $\{1,5,10,15\}$.
 In the final experiment, we compare the recovery performance of AGD for different numbers of IRSs $D$.
-As shown in Figure 3, the recovery error increases linearly with $D$, aligning with the theoretical analysis in Theorem 5.
-Additionally, as $T$ increases, the recovery error decreases.
-Thus, for fixed $T$ time slots, a trade-off emerges between mitigating propagation attenuations with multiple IRSs and minimizing recovery error.
-VI Conclusion
-This paper addresses the significant gap in theoretical error analysis concerning channel estimation for IRS-assisted MIMO systems.
-By establishing the equivalence between these systems and tensor train-based tensor-on-tensor (ToT) regression, we provide insights into the fundamental factors crucial for the successful recovery of channel matrices.
-Our analysis highlights the pivotal role of the relationship between the number of user terminals or base stations and the number of time slots for ensuring stable recovery.
-Furthermore, we extend our investigation to consider low-rank channel matrices and unknown IRS, enhancing the applicability of our findings.
+As shown in Figure 5, the recovery error increases linearly with $D$, aligning with the theoretical analysis in Theorems 5 and 6.
+Moreover, for a fixed number of time slots $T$, the results highlight a fundamental trade-off between enhancing signal propagation through additional IRS layers and maintaining low recovery error.
+Note that while increasing $D$ can improve end-to-end channel coverage, it simultaneously escalates the difficulty of accurate channel estimation due to the higher dimensionality and the accumulation of noise.
+V Conclusion
+This paper addresses the significant gap in optimal error analysis concerning channel estimation for IRS-assisted MIMO systems.
+By establishing the equivalence between these systems and tensor train-based ToT regression, we provide insights into the fundamental factors crucial for the successful recovery of channel matrices.
+Specifically, we have established a theoretical upper bound on the recovery error for the constrained least-squares solution, along with a minimax lower bound for the IRS-assisted MIMO model.
+Our results demonstrate that the recovery error decreases inversely with the number of time slots and increases proportionally with the number of unknown entries in the channel matrices.
+These findings quantify the fundamental trade-offs inherent in IRS-assisted MIMO channel estimation and provide theoretical justification for the observed empirical behaviors.
 Additionally, through our exploration of a multi-hop IRS scheme, we evaluate corresponding recovery errors, shedding light on the performance of such configurations.
-Finally, to validate our theoretical conclusions, we propose and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–demonstrating their effectiveness through experimental results.
+Finally, we validate our theoretical findings through extensive simulations, which consistently align with the derived recovery bounds.
+Building upon the results established in this paper, several important research directions can be identified for further investigation.
   • An important area for future research involves conducting error analysis on two-way, two-hop IRS-assisted MIMO systems [87, 88, 89].
@@
 Thus, our current analytical approach cannot be directly extended to this setup.
-Furthermore, another challenging aspect for future investigation is the local convergence analysis and global geometry of the IRS-assisted optimization problem.
-Despite the promising experimental results obtained even with random initialization, unlike the rotation ambiguity commonly encountered in standard nonconvex matrix/tensor recovery [45, 46, 47, 48], a diagonal (scaling) ambiguity exists between channel matrices, as noted in [36, Section 2].
+  • Another important challenge is extending our results to the downlink communication scenario.
+Notably, our recovery guarantees remain applicable in the downlink setting, provided that the RIP holds–achieved by simply interchanging the roles of transmitter and receiver.
+However, as evidenced by prior studies [90, 91, 92, 93], the number of time slots in the downlink can fall below the number of BSs, thereby violating the RIP condition.
+This limitation underscores the necessity of developing new analytical frameworks tailored to the downlink regime.
+  • Furthermore, recent studies have revealed that even when the direct BS–UT link is blocked, the scattering matrix of the direct channel remains nonzero [94].
+Under specific assumptions, the IRS-assisted MIMO model can therefore be reformulated as in [95, eq.
+(80)], which explicitly incorporates the scattering matrix of the reconfigurable impedance network and the IRS reflection (or scattering) coefficient matrix.
+When these matrices are known, our theoretical results remain directly applicable.
+However, in practical scenarios, the scattering matrix is typically unknown and the model includes an inverse operation, which prevents the direct application of our current analytical approach.
+Consequently, extending the error analysis to models without explicit scattering information remains an open and important direction for future work.
+  • Finally, a further avenue for exploration is the local convergence analysis and global geometry of the IRS-assisted optimization problem.
+Despite the promising experimental results obtained even with random initialization, unlike the rotation ambiguity commonly encountered in standard nonconvex matrix/tensor recovery [47, 48, 49, 51], a diagonal (scaling) ambiguity exists between channel matrices, as noted in [37, Section 2].
 This ambiguity poses a challenge in analyzing the optimization problem at hand.
@@
 Thus, the TT format can be represented by $D$ tensor factors $\{{\bm{B}}_{d}\}_{d\geq 1}$, with a total of $O(D\overline{N}\overline{r}^{2})$ parameters, where $\overline{N}=\max_{d}N_{d}$ and $\overline{r}=\max_{d}r_{d}$.
-In addition, for any two TT format tensors ~,^$\widetilde{\mathcal{B}},\widehat{\mathcal{B}}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ with factors $\{\widetilde{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widetilde{r}_{d-1}\times\widetilde{r}_{d}}\}$ and $\{\widehat{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widehat{r}_{d-1}\times\widehat{r}_{d}}\}$, each element of the summation $\mathcal{B}=\widetilde{\mathcal{B}}+\widehat{\mathcal{B}}$ can be represented by
-$\mathcal{B}(s_{1},\dots,s_{D})=\begin{bmatrix}\widetilde{\bm{B}}_{1}(s_{1})\!\!\!\!&\widehat{\bm{B}}_{1}(s_{1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{2}(s_{2})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{2}(s_{2})\end{bmatrix}$
-$\cdots\begin{bmatrix}\widetilde{\bm{B}}_{D-1}(s_{D-1})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{D-1}(s_{D-1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{D}(s_{D})\\ \widehat{\bm{B}}_{D}(s_{D})\end{bmatrix},$ (30)
+In addition, for any two TT format tensors $\widetilde{\mathcal{B}},\widehat{\mathcal{B}}\in\mathbb{C}^{N_{1}\times\cdots\times N_{D}}$ with factors $\{\widetilde{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widetilde{r}_{d-1}\times\widetilde{r}_{d}}\}$ and $\{\widehat{\bm{B}}_{d}(s_{d})\in\mathbb{C}^{\widehat{r}_{d-1}\times\widehat{r}_{d}}\}$, each element of the summation $\mathcal{B}=\widetilde{\mathcal{B}}$$\widehat{\mathcal{B}}$ can be represented by
+$\hskip 0.0pt\mathcal{B}(s_{1},\dots,s_{D})$ $\!\!\!\!=\!\!\!\!$ $\begin{bmatrix}\widetilde{\bm{B}}_{1}(s_{1})\!\!\!\!&\widehat{\bm{B}}_{1}(s_{1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{2}(s_{2})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{2}(s_{2})\end{bmatrix}\cdots$ (31)
+$\hskip-28.45274pt\begin{bmatrix}\widetilde{\bm{B}}_{D-1}(s_{D-1})\!\!\!\!&{\bm{0}}\\ {\bm{0}}\!\!\!\!&\widehat{\bm{B}}_{D-1}(s_{D-1})\end{bmatrix}\begin{bmatrix}\widetilde{\bm{B}}_{D}(s_{D})\\ \widehat{\bm{B}}_{D}(s_{D})\end{bmatrix}$,
 which implies that $\mathcal{B}$ can also be represented in the TT format with ranks $r_{d}\leq\widetilde{r}_{d}+\widehat{r}_{d}$ for $d=1,\dots,D-1$.
-Since any tensor can be decomposed in the TT format (29) with sufficiently large TT ranks [37, Theorem 2.1], the decomposition of a tensor $\mathcal{B}$ into the form (29) is generally not unique: not only are the factors ${\bm{B}_{d}(:,s_{d},:)}$ not unique, but also the dimensions of these factors can vary.
+Since any tensor can be decomposed in the TT format (30) with sufficiently large TT ranks [38, Theorem 2.1], the decomposition of a tensor $\mathcal{B}$ into the form (30) is generally not unique: not only are the factors $\bm{B}_{d}(:,s_{d},:)$} not unique, but also the dimensions of these factors can vary.
 To introduce the factorization with the smallest possible dimensions $\bm{r}=(r_{1},\ldots,r_{D-1})$, for convenience, for each $d$, we put $\{{\bm{B}_{d}(:,s_{d},:)}\}_{s_{d}=1}^{N_{d}}$ together into the following two forms
-$L(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)\\ \vdots\\ \bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{(r_{d-1}N_{d})\times r_{d}},$ (31)
-$R(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)&\cdots&\bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{r_{d-1}\times(N_{d}r_{d})},$ (32)
+$\hskip-28.45274ptL(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)\\ \vdots\\ \bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{(r_{d-1}N_{d})\times r_{d}},$ (32)
+$\hskip-28.45274ptR(\bm{B}_{d})=\begin{bmatrix}\bm{B}_{d}(1)&\cdots&\bm{B}_{d}(N_{d})\end{bmatrix}\in\mathbb{C}^{r_{d-1}\times(N_{d}r_{d})},$ (33)
 where $L(\bm{B}_{d})$ and $R(\bm{B}_{d})$ are often called the left and right unfoldings of $\bm{B}_{d}$, respectively, if we view $\bm{B}_{d}$ as a tensor.
@@
 According to [96], there is exactly one set of ranks $\bm{r}$ that $\mathcal{B}$ admits a minimal TT decomposition.
-Appendix B Proof of Theorem 1
-Proof.
-By substituting the concentration inequality from [64, Theorem 2.3] with the concentration inequality for complex-valued subgaussian random variables, when (9) holds true, we can assert with a probability of $1-e^{-cT}$ that:
-[…]
-[…]
-Due to $\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\sum_{p,k}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ and $\|\mathcal{B}\|_{F}^{2}=\sum_{p,k}\|\mathcal{B}(p,k,:)\|_{F}^{2}$, this completes the proof. ∎
+Appendix B Nonconvex Iterative Optimization Algorithms
+In this section, we restate the nonconvex iterative algorithms used to estimate the channel matrices within the framework of the TT-based ToT regression model.
+For a single-hop IRS-assisted MIMO system, we directly adopt the bilinear alternating least squares (BALS) algorithm [31, Algorithm 2].
+The computational complexity of each iteration of the BALS method, measured in terms of multiplication operations, is given by $O\big(TUMN+TLPKN+KTN\cdot\min\{KT,N\},LPKN\cdot\min\{LPK,N\}\big)$.
+In contrast, for a multi-hop IRS-assisted MIMO system, a closed-form solution for the alternating least squares method is not tractable.
+Therefore, we primarily consider the alternating gradient descent (AGD) method proposed in [77].
+Note that we ignore the direct transmission path and focus on minimizing the following loss function:
+$\operatorname*{\text{arg~min}}_{\widehat{\bm{B}}_{D}\in\mathbb{C}^{LP\times N_{D}},\widehat{\bm{B}}_{0}\in\mathbb{C}^{N_{1}\times UM},\atop\widehat{\bm{B}}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]}g(\widehat{\bm{B}}_{D},\dots,\widehat{\bm{B}}_{0})$ (34)
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}([\widehat{\bm{B}}_{D},\mathcal{S}_{D},\dots,\widehat{\bm{B}}_{1},\mathcal{S}_{1},\widehat{\bm{B}}_{0}])-\widetilde{\mathcal{Y}}\|_{F}^{2}.$
+We solve this nonconvex optimization problem by the following AGD algorithm for $d=0,\dots,D$:
+$\bm{B}_{d}^{(l+1)}=\bm{B}_{d}^{(l)}\!\!-\!\mu\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D}^{(l+1)},\dots,\bm{B}_{d+1}^{(l+1)},\bm{B}_{d}^{(l)},\dots,\bm{B}_{0}^{(l)}),$
+where $\mu$ is a step size and the Wirtinger gradient $\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})$ is defined as:
+$\hskip-14.22636pt\nabla_{\bm{B}_{d}^{*}}g(\bm{B}_{D},\dots,\bm{B}_{0})=\frac{1}{T}\sum_{k=1}^{K}\mathcal{S}_{d+1}^{\rm H}(:,k,:)\bm{B}_{d+1}^{\rm H}\cdots\mathcal{S}_{D}^{\rm H}(:,k,:)$
+$\hskip 5.69046pt\cdot\bm{B}_{D}^{\rm H}\bigg(\mathcal{X}([\bm{B}_{D},\mathcal{S}_{D}(:,k,:),\dots,\bm{B}_{1},\mathcal{S}_{1}(:,k,:),\bm{B}_{0}])$
+$\hskip 5.69046pt-\widetilde{\mathcal{Y}}(:,k,:)\bigg)\bm{X}^{\rm H}\bm{B}_{0}^{\rm H}\cdots\bm{B}_{d-1}^{\rm H}\mathcal{S}_{d}^{\rm H}(:,k,:).$ (35)
+It is worth noting that AGD in this scenario operates without coefficient constraints, with the only requirement being $T=\Omega(UM)$ to ensure the RIP condition holds.
+Defining $N=\max_{d}N_{d}$, the overall computational complexity of each AGD iteration, measured in terms of multiplication operations, is $O\big(KD(DN^{3}+(N+T)LP+UMT(N+LP))\big)$.
 Appendix C Proof of Theorem 1
 Proof.
-Using (7), we have
-[…]
-[…]
+By substituting the concentration inequality from [65, Theorem 2.3] with the concentration inequality for complex-valued subgaussian random variables, when (9) in the main paper holds true, we can assert with a probability of $1-e^{-cT}$ that:
+$(1-\delta_{UM})\|\mathcal{B}(p,k,:)\|_{F}^{2}$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ (36)
+$\!\!\!\!\leq\!\!\!\!$ $(1+\delta_{UM})\|\mathcal{B}(p,k,:)\|_{F}^{2}.$
+Due to $\|\mathcal{X}(\mathcal{B})\|_{F}^{2}=\sum_{p,k}\|\mathcal{X}(\mathcal{B})(p,k,:)\|_{2}^{2}$ and $\|\mathcal{B}\|_{F}^{2}=\sum_{p,k}\|\mathcal{B}(p,k,:)\|_{F}^{2}$, this completes the proof for complex-valued subgaussian random variables.
+In addition, it is straightforward to see that when $T\geq UM$, the normalized DFT matrix satisfies the RIP. ∎
+Appendix D Proof of Theorem 2
+Proof.
+Using (7) in the main paper, we have
+$0$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})-\mathcal{Y}\|_{F}^{2}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})-\mathcal{Y}\|_{F}^{2}$ (37)
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})-\mathcal{X}(\mathcal{B}^{\star})-\mathcal{W}\|_{F}^{2}$
 $-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})-\mathcal{X}(\mathcal{B}^{\star})-\mathcal{W}\|_{F}^{2}$
-[…]
+$\!\!\!\!=\!\!\!\!$ $\frac{2}{T}\operatorname{Re}\left\{\langle\mathcal{X}(\mathcal{B}^{\star})+\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle\right\}$
 $+\frac{1}{T}\|\mathcal{X}(\mathcal{B}^{\star})\|_{F}^{2}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}})\|_{F}^{2}$
-[…]
+$\!\!\!\!=\!\!\!\!$ $\frac{2}{T}\operatorname{Re}\left\{\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle\right\}-\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2},$
 which further implies that
 $\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2}\leq\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|.$ (38)
-According to (11), we can directly obtain
+According to (11) in the main paper, we can directly obtain
 $\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\|_{F}^{2}\geq(1-\delta_{UM})\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}^{2}.$ (39)
@@
 $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$ (40)
-[…]
-[…]
+$\!\!\!\!=\!\!\!\!$ $\frac{2\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N,\mathcal{S}},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$\!\!\!\!=\!\!\!\!$ $\frac{2\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}}{T}\hskip-14.22636pt\max_{\|\bm{H}_{1}\|\leq 1,\|\bm{H}_{2}\|_{F}\leq 1,\atop\|L(\widetilde{\mathcal{S}})\|\leq 1}\hskip-21.33955pt|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
 where $\bm{X}^{*}$ is the conjugate matrix of $\bm{X}$ and $\mathcal{W}\times_{3}^{2}\bm{X}^{*}=\sum_{t=1}^{T}\mathcal{W}(:,:,t)\bm{X}^{*}(:,t)$.
-According to [48, eq.(44)], the last line follows $\|\mathcal{H}\|_{F}=\|\bm{H}_{2}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
+According to [51, eq.(44)], the last line follows $\|\mathcal{H}\|_{F}=\|\bm{H}_{2}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A of the main paper.
 To begin, according to [97], we can construct an $\epsilon$-net $\{\bm{H}_{1}^{(1)},\dots,\bm{H}_{1}^{(n_{1})}\}$ with the covering number $n_{1}\leq(\frac{4+\epsilon}{\epsilon})^{LPN}$ for the set of factors $\{\bm{H}_{1}\in\mathbb{C}^{LP\times N}:\|\bm{H}_{1}\|\leq 1\}$ such that
@@
 $\sup_{\bm{H}_{2}:\|\bm{H}_{2}\|_{F}\leq 1}\min_{p_{2}\leq n_{2}}\|\bm{H}_{2}-\bm{H}_{2}^{(p_{2})})\|_{F}\leq\epsilon.$ (42)
-Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number
-$n_{1}n_{2}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{LPN+UMN}$ (40)
-for any TT format tensor $\mathcal{H}=[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}]\in\mathbb{B}_{2N,\widetilde{\mathcal{S}}}$.
+Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number $n_{1}n_{2}\leq(\frac{4+\epsilon}{\epsilon})^{LPN+UMN}$ for any TT format tensor $\mathcal{H}=[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}]\in\mathbb{B}_{2N,\widetilde{\mathcal{S}}}$.
 Denote by $A$ the value of (40), i.e.,
-$[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]=\operatorname*{\text{arg~{}max}}_{\|\bm{H}_{1}\|\leq 1,\|\bm{H}_{2}\|_{F}\leq 1,\atop\|L(\widetilde{\mathcal{S}})\|\leq 1}1/T|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$ (41)
-$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (42)
-Using $\mathcal{I}$ to denote the index set $[n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
+$\hskip-34.14322pt[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]=\hskip-14.22636pt\operatorname*{\text{arg~max}}_{\|\bm{H}_{1}\|\leq 1,\atop\|\bm{H}_{2}\|_{F}\leq 1,\|L(\widetilde{\mathcal{S}})\|\leq 1}\hskip-11.38092pt\frac{1}{T}|\langle[\bm{H}_{1},\widetilde{\mathcal{S}},\bm{H}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$ (43)
+$\hskip-34.14322ptA:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (44)
+Using $\mathcal{I}$ to denote the index set $n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
 $\|\widetilde{\bm{H}}_{1}-\bm{H}_{1}^{(p_{1})}\|\leq\epsilon,\ \ \ \text{and}\ \ \ \|\widetilde{\bm{H}}_{2}-\bm{H}_{2}^{(p_{2})}\|_{F}\leq\epsilon,$ (45)
 and taking $\epsilon=\frac{1}{4}$ gives
-[…]
+$A$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$ (46)
 $+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]-[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
-[…]
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
 $+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{1}-\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}]$
 $+[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\widetilde{\bm{H}}_{2}-\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
-[…]
-[…]
+$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}^{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+2\epsilon A$
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}^{(p_{1})},\widetilde{\mathcal{S}},\bm{H}_{2}{(p_{2})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{A}{2}.$
 Note that each element in $\mathcal{W}$ follows the complex normal distribution $\mathcal{C}\mathcal{N}(0,\gamma^{2})$.
@@
 $\operatorname{\mathbb{P}}\left(\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq t|\bm{X}\right)\leq e^{-\frac{T^{2}t^{2}}{2\gamma^{2}\|\mathcal{X}(\mathcal{H}^{(p)})\|_{F}^{2}}}.$ (47)
-[…]
+Furthermore, under the event $F:=\{\mathcal{X}\text{ satisfies $UM$-RIP}$  $\text{ with constant $\delta_{UM}$}\}$, which implies that $\frac{1}{T}\|\mathcal{X}(\mathcal{H}^{(p)})\|_{F}^{2}\leq(1+\delta_{UM})\|\mathcal{H}^{(p)}\|_{F}^{2}$.
+Plugging this together with the fact $\|\mathcal{H}^{(p)}\|_{F}\leq 1$ into the above further gives
 $\operatorname{\mathbb{P}}\left(\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq t|F\right)\leq e^{-\frac{Tt^{2}}{2(1+\delta_{UM})\gamma^{2}}}.$ (48)
 We now apply this tail bound to (46) and get
-[…]
-[…]
-[…]
+$\operatorname{\mathbb{P}}\left(A\geq t|F\right)$ $\!\!\!\!\leq\!\!\!\!$ $\operatorname{\mathbb{P}}\left(\max_{p_{1},p_{2}}\frac{1}{T}|\langle\mathcal{H}^{(p)},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|\geq\frac{t}{2}|F\right)$ (49)
+$\!\!\!\!\leq\!\!\!\!$ $\big(\frac{4+\epsilon}{\epsilon}\big)^{2LPN+2UMN}e^{-\frac{Tt^{2}}{8(1+\delta_{UM})\gamma^{2}}}$
+$\!\!\!\!\leq\!\!\!\!$ $e^{-\frac{Tt^{2}}{8(1+\delta_{UM})\gamma^{2}}+c_{1}(LPN+UMN)},$
 where $c_{1}$ is a constant and based on the assumption $\epsilon=\frac{1}{4}$ in (46), $\frac{4+\epsilon}{\epsilon}=17$.
@@
 $\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma\right)$ (50)
-[…]
-[…]
-[…]
-[…]
-where $c_{i},i=3,4,5$ are constants. Note that $P(F)$ follows Theorem 1.
+$\!\!\!\!\geq\!\!\!\!$ $\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma\cap F\right)$
+$\!\!\!\!\geq\!\!\!\!$ $P(F)\operatorname{\mathbb{P}}\left(A\leq\frac{c_{2}\sqrt{(1+\delta_{UM})(LPN+UMN)}}{\sqrt{T}}\gamma|F\right)$
+$\!\!\!\!\geq\!\!\!\!$ $(1-e^{-c_{3}UM})(1-e^{-c_{4}(LPN+UMN)})$
+$\!\!\!\!\geq\!\!\!\!$ $1-2e^{-c_{5}(LPN+UMN)},$
+where $c_{i},i=3,4,5$ are constants. Note that $P(F)$ follows Theorem 1 in the main paper.
 Combing (39), we can obtain
-^$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\leq O\bigg{(}\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg{)}.$ (49)
-Appendix D Proof of Theorem 3
+$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(LPN+UMN)}}{(1-\delta_{UM})\sqrt{T}}\bigg).$ (51)
+Appendix E Proof of (II-C) in the main paper
+Proof.
+Similar to the analysis presented in Appendix D, we focus on deriving an upper bound for the term $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$.
+To this end, we first rewrite the term as follows:
+$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}-\mathcal{B}^{\star})\rangle|$ $\!\!\!\!=\!\!\!\!$ $2/T\hskip-22.76228pt\max_{\mathcal{B}\in\{\mathcal{C}:\mathcal{C}\in\mathbb{C}^{LP\times K\times UM},\atop\|\mathcal{C}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\}}\hskip-22.76228pt|\langle\mathcal{B},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (52)
+Without loss of generality, we assume $\widehat{\mathcal{B}},\mathcal{B}^{\star}\in\mathbb{B}=\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}},\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}},\|\bm{\Sigma}_{1}\|_{F}\leq L_{1},\|\bm{\Sigma}_{2}\|_{F}\leq L_{2},\|\bm{\Sigma}_{1}\|_{0}=s_{1},\|\bm{\Sigma}_{2}\|_{0}=s_{2}\ \text{are unknown and}\ \mathcal{S}\in\mathbb{C}^{N\times K\times N},\bm{H}_{1}\in\mathbb{C}^{LP\times F_{1}},\bm{H}_{2}\in\mathbb{C}^{F_{1}\times N},\bm{G}_{1}\in\mathbb{C}^{N\times F_{2}}\ \text{and}\ \bm{G}_{2}\in\mathbb{C}^{F_{2}\times UM}\ \text{are known}\}$.
+According to [98, eq.(2)], we can construct an $\epsilon_{1}$-net $\{\bm{\Sigma}_{1}^{(1)},\dots,\bm{\Sigma}_{1}^{(n_{1})}\}$ with the covering number $n_{1}\leq(\frac{CF_{1}L_{1}}{s_{1}\epsilon_{1}})^{s_{1}}$ where $C$ is a positive constant, for the set of factors $\{\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}:\|\bm{\Sigma}_{1}\|_{F}\leq L_{1}\ \text{and}\ \|\bm{\Sigma}_{1}\|_{0}=s_{1}\}$ such that
+$\sup_{\bm{\Sigma}_{1}\in\mathbb{C}^{F_{1}\times F_{1}}:\|\bm{\Sigma}_{1}\|_{F}\leq L_{1}\atop\ \text{and}\ \|\bm{\Sigma}_{1}\|_{0}=s_{1}}\min_{p_{1}\leq n_{1}}\|\bm{\Sigma}_{1}-\bm{\Sigma}_{1}^{(p_{1})}\|\leq\epsilon_{1}.$ (53)
+Similarly, we can construct an $\epsilon_{2}$-net $\{\bm{\Sigma}_{2}^{(1)},\dots,\bm{\Sigma}_{2}^{(n_{2})}\}$ with the covering number $n_{2}\leq(\frac{CF_{2}L_{2}}{s_{2}\epsilon_{2}})^{s_{2}}$ for the set of factors $\{\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}:\|\bm{\Sigma}_{2}\|_{F}\leq L_{2}\ \text{and}\ \|\bm{\Sigma}_{2}\|_{0}=s_{2}\}$ such that
+$\sup_{\bm{\Sigma}_{2}\in\mathbb{C}^{F_{2}\times F_{2}}:\|\bm{\Sigma}_{2}\|_{F}\leq L_{2}\atop\ \text{and}\ \|\bm{\Sigma}_{2}\|_{0}=s_{2}}\min_{p_{2}\leq n_{2}}\|\bm{\Sigma}_{2}-\bm{\Sigma}_{2}^{(p_{2})}\|\leq\epsilon_{2}.$ (54)
+Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{1}n_{2})}\}$ with covering number $n_{1}n_{2}\!\!\leq\!\!(\frac{CF_{1}L_{1}}{s_{1}\epsilon_{1}})^{s_{1}}(\frac{CF_{2}L_{2}}{s_{2}\epsilon_{2}})^{s_{2}}$ for any TT format tensor $\mathcal{H}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]$.
+Denote by $B$ the value of (52), i.e.,
+$\hskip-22.76228pt[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}]=\hskip-14.22636pt\operatorname*{\text{arg~max}}_{\mathcal{B}\in\{\mathcal{C}:\mathcal{C}\in\mathbb{C}^{LP\times K\times UM},\atop\|\mathcal{C}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}\}}\frac{1}{T}|\langle\mathcal{B},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
+$\hskip-22.76228ptB:=\frac{1}{T}|\langle[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$
+Using $\mathcal{I}$ to denote the index set $n_{1}]\times[n_{2}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{1},p_{2})\in\mathcal{I}$ such that
+$\|\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}\|\leq\epsilon_{1},\ \ \ \text{and}\ \ \ \|\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}\|\leq\epsilon_{2}.$ (55)
+Then taking $\epsilon_{1}=\frac{L_{1}}{4}$ and $\epsilon_{2}=\frac{L_{2}}{4}$ gives
+$B$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$+\frac{1}{T}|\langle[\bm{H}_{1}\widehat{\bm{\Sigma}}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}]$
+$-[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$+\frac{\epsilon_{1}}{T}|\langle[\bm{H}_{1}\frac{(\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}^{(p_{1})})}{\|\widehat{\bm{\Sigma}}_{1}-\bm{\Sigma}_{1}^{(p_{1})}\|_{F}}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\widehat{\bm{\Sigma}}_{2}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$+\frac{\epsilon_{2}}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\frac{(\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}^{(p_{2})})}{\|\widehat{\bm{\Sigma}}_{2}-\bm{\Sigma}_{2}^{(p_{2})}\|_{F}}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{B}{2}.$
+We further assume that $\bm{H}_{1}\bm{\Sigma}_{1}^{(p_{1})}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}^{(p_{2})}\bm{G}_{2}]\in\mathbb{B}\cap\{\mathcal{B}=[\bm{H}_{1}\bm{\Sigma}_{1}\bm{H}_{2},\mathcal{S},\bm{G}_{1}\bm{\Sigma}_{2}\bm{G}_{2}]:\|\mathcal{B}\|_{F}\leq\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}\}$.
+Following the same analysis of (51), with high probability, we have
+$\|\widehat{\mathcal{B}}-\mathcal{B}^{\star}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(s_{1}\log(\frac{F_{1}}{s_{1}})+s_{2}\log(\frac{F_{2}}{s_{2}}))}}{(1-\delta_{UM})\sqrt{T}}\bigg).$
+Appendix F Proof of Theorem 3
 Proof.
 Suppose we can find a set of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that $\min_{j\neq k}\|\mathcal{B}^{j}-\mathcal{B}^{k}\|_{F}\geq s$.
-According to [84, Theorem 4], when each element of $\bm{X}$ and $\mathcal{W}$ respectively follow $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, we have
+According to [72, Theorem 4], when each element of $\bm{X}$ and $\mathcal{W}$ respectively follow $\mathcal{C}\mathcal{N}(0,1)$ and $\mathcal{C}\mathcal{N}(0,\gamma^{2})$, we have
 $\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}$ (56)
-[…]
+$\!\!\!\!\geq\!\!\!\!$ $\frac{s}{2}\bigg(1-\frac{\frac{T}{2\gamma^{2}}\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}^{2}+\log 2}{\log n}\bigg).$
 Next, we consider one construction for the sets of $\{\mathcal{B}^{j}\}_{j=1}^{n}\in\mathbb{B}_{N,\mathcal{S}}$ such that we can obtain a proper lower bound for $\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$ and a proper upper bound for $\max_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}$.
 Given the equivalence between any two tensors $\mathcal{B}^{j_{1}}=[\bm{A}_{1}^{j_{1}},\mathcal{S},\bm{A}_{2}]$, $\mathcal{B}^{j_{2}}=[\bm{A}_{1}^{j_{2}},\mathcal{S},\bm{A}_{2}]$ and their canonical forms [51], the TT-SVD can be applied to transform these tensors into canonical formats $\mathcal{B}^{j_{1}}=[\widetilde{\bm{A}}_{1}^{j_{1}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ and $\mathcal{B}^{j_{2}}=[\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]$ with $R(\widetilde{\mathcal{S}})R^{\rm H}(\widetilde{\mathcal{S}})={\bf I}_{N}$ and $\widetilde{\bm{A}}_{2}\widetilde{\bm{A}}_{2}^{\rm H}={\bf I}_{N}$.
-Given any $\delta>0$ and $N\geq C^{\prime}$ by [85, Lemma 5], we can construct a set of $\{\widetilde{\bm{A}}_{1}^{1},\dots,\widetilde{\bm{A}}_{1}^{n_{0}}\}$ with cardinality $n_{0}\geq\frac{1}{4}e^{\frac{LPN}{128}}$ such that: (1) $\|\widetilde{\bm{A}}_{1}^{j}\|_{F}=\delta$ holds for all $j=1,\dots,n_{0}$, (2) $\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta$ for all $j_{1},j_{2}\in[n_{0}],j_{1}\neq j_{2}$.
+Given any $\delta>0$ and $N\geq C^{\prime}$ by [99, Lemma 5], we can construct a set of $\{\widetilde{\bm{A}}_{1}^{1},\dots,\widetilde{\bm{A}}_{1}^{n_{0}}\}$ with cardinality $n_{0}\geq\frac{1}{4}e^{\frac{LPN}{128}}$ such that: (1) $\|\widetilde{\bm{A}}_{1}^{j}\|_{F}=\delta$ holds for all $j=1,\dots,n_{0}$, (2) $\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta$ for all $j_{1},j_{2}\in[n_{0}],j_{1}\neq j_{2}$.
 Due to $\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}=\|[\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}},\widetilde{\mathcal{S}},\widetilde{\bm{A}}_{2}]\|_{F}=\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}$, we can further get
@@
 $\min_{j_{1}\neq j_{2}}\|\mathcal{B}^{j_{1}}-\mathcal{B}^{j_{2}}\|_{F}=\min_{j_{1}\neq j_{2}}\|\widetilde{\bm{A}}_{1}^{j_{1}}-\widetilde{\bm{A}}_{1}^{j_{2}}\|_{F}\geq\delta.$ (57)
-Then we plug (D) into (50) and have
-[…]
-[…]
+Then we plug (F) into (56) and have
+$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}$ $\!\!\!\!\geq\!\!\!\!$ $\frac{\delta}{2}\bigg(1-\frac{\frac{2T\delta^{2}}{\gamma^{2}}+\log 2}{c_{1}LPN}\bigg)$ (58)
+$\!\!\!\!\geq\!\!\!\!$ $c_{2}\sqrt{\frac{LPN}{T}}\gamma,$
 where the second inequality with a constant $c_{2}>0$ follows $\delta=c_{3}\sqrt{\frac{LPN}{m}}\gamma$ for a constant $c_{3}>0$.
@@
 Then following the previous analysis, we can get
-$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{UMN}{T}}\gamma\bigg{)}.$ (53)
+$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}=\Omega\bigg(\sqrt{\frac{UMN}{T}}\gamma\bigg).$ (59)
 Finally, we sum (58)-(59) and calculate the average to derive
-$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{LPN+UMN}{2T}}\gamma\bigg{)}.$ (54)
-Appendix E Proof of Theorem 5
+$\inf_{\widehat{\mathcal{B}}}\sup_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\operatorname{\mathbb{E}}{\|\widehat{\mathcal{B}}-\mathcal{B}\|_{F}}=\Omega\bigg(\sqrt{\frac{LPN+UMN}{2T}}\gamma\bigg).$ (60)
+Appendix G Proof of Theorem 5
 Proof.
@@
 $\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\|_{F}^{2}\leq\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|.$ (61)
-First, according to Theorem 4, we have
+First, according to Theorem 4 in the main paper, we have
 $\frac{1}{T}\|\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\|_{F}^{2}\geq(1-\delta_{UM})\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}^{2}.$ (62)
 Additionally, we can also rewrite $\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$ as follows:
-$\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$
-$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N}^{D},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
-$=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (57)
-The last line follows that $\|\mathcal{H}\|_{F}^{2}=\|\bm{H}_{0}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A.
+$\hskip-14.22636pt\frac{2}{T}|\langle\mathcal{W},\mathcal{X}(\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D})\rangle|$
+$\hskip-14.22636pt=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\max_{\mathcal{H}\in\mathbb{B}_{2N}^{D},\|\mathcal{H}\|_{F}\leq 1}|\langle\mathcal{H},\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
+$\hskip-14.22636pt=\frac{2\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}}{T}\hskip-31.2982pt\max_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}\hskip-36.98866pt|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (63)
+The last line follows that $\|\mathcal{H}\|_{F}^{2}=\|\bm{H}_{0}\|_{F}\leq 1$ for a left-orthogonal TT form in Appendix A of the main paper.
 Next, we can construct an $\epsilon$-net $\{\bm{H}_{i}^{(1)},\dots,\bm{H}_{i}^{(n_{i})}\}$ with the covering number $n_{i}\leq(\frac{4+\epsilon}{\epsilon})^{N_{i}N_{i+1}},i\in[D]$ for the set of factors $\{\bm{H}_{i}\in\mathbb{C}^{N_{i+1}\times N_{i}}:\|\bm{H}_{i}\|\leq 1\}$ such that
@@
 $\sup_{\bm{H}_{0}:\|\bm{H}_{0}\|_{F}\leq 1}\min_{p_{0}\leq n_{0}}\|\bm{H}_{0}-\bm{H}_{0}^{(p_{0})}\|_{F}\leq\epsilon.$ (65)
-Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{0}\cdots n_{D})}\}$ with covering number
-$n_{0}\cdots n_{D}\leq\bigg{(}\frac{4+\epsilon}{\epsilon}\bigg{)}^{\sum_{d=0}^{D}N_{d}N_{d+1}}$ (60)
-for any $\mathcal{H}\in\mathbb{B}_{2N,\{\widetilde{\mathcal{S}}_{i}\}}^{D}$.
-Denote by $A$ the value of (E), i.e.,
+Therefore, we can construct an $\epsilon$-net $\{\mathcal{H}^{(1)},\ldots,\mathcal{H}^{(n_{0}\cdots n_{D})}\}$ with covering number $n_{0}\cdots n_{D}\leq(\frac{4+\epsilon}{\epsilon})^{\sum_{d=0}^{D}N_{d}N_{d+1}}$ for any $\mathcal{H}\in\mathbb{B}_{2N,\{\widetilde{\mathcal{S}}_{i}\}}^{D}$.
+Denote by $A$ the value of (G), i.e.,
 $[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]$ (66)
-[…]
-$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$
-Using $\mathcal{I}$ to denote the index set $[n_{0}]\times\cdots\times[n_{D}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{0},\dots,p_{D})\in\mathcal{I}$ such that
+$\!\!\!\!\hskip-14.22636pt=\!\!\!\!$ $\hskip-19.91684pt\operatorname*{\text{arg~max}}_{\|\bm{H}_{i}\|\leq 1,i\in[D],\atop\|L(\widetilde{\mathcal{S}}_{i})\|\leq 1,i\in[D],\|\bm{H}_{0}\|_{F}\leq 1}\hskip-25.6073pt|\langle[\bm{H}_{D},\widetilde{\mathcal{S}}_{D},\dots,\bm{H}_{1},\widetilde{\mathcal{S}}_{1},\bm{H}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|,$
+$A:=\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\bm{H}}_{1},\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|.$ (67)
+Using $\mathcal{I}$ to denote the index set $n_{0}]\times\cdots\times[n_{D}]$, then according to the construction of the $\epsilon$-net, there exists $p=(p_{0},\dots,p_{D})\in\mathcal{I}$ such that
 $\|\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})}\|\leq\epsilon,i\in[D]\ \text{and}\ \|\widetilde{\bm{H}}_{0}-\bm{H}_{0}^{(p_{0})}\|_{F}\leq\epsilon,$ (68)
 and taking $\epsilon=\frac{1}{2(D+1)}$ gives
-[…]
+$A$ $\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$ (69)
 $+\frac{1}{T}|\langle[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]$
 $-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
-[…]
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
 $+\frac{1}{T}\bigg|\sum_{i=0}^{D}\bigg\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\bm{H}_{D-1}^{(p_{D-1})},\dots,\bm{H}_{i+1}^{(p_{D_{i+1}})},\widetilde{\mathcal{S}}_{i+1},$
 $\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})},\widetilde{\mathcal{S}}_{i},\widetilde{\bm{H}}_{i-1},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\bigg\rangle\bigg|$
-[…]
+$\!\!\!\!\leq\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|$
 $+(D+1)\epsilon A$
-[…]
-where the first equation follows the fact that $(l,k,u)$-th element of $[\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}]$ can be decomposed into $\sum_{i=0}^{D}\bm{H}_{D}^{(p_{D})}(l,:)\widetilde{\mathcal{S}}_{D}(:,k,:)\bm{H}_{D-1}^{(p_{D-1})}\\ \cdots\bm{H}_{i+1}^{(p_{D_{i+1}})}\widetilde{\mathcal{S}}_{i+1}(:,k,:)(\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})})\widetilde{\mathcal{S}}_{i}(:,k,:)\widetilde{\bm{H}}_{i-1}\widetilde{\mathcal{S}}_{1}(:,k,:)\widetilde{\bm{H}}_{0}(:,u)$.
+$\!\!\!\!=\!\!\!\!$ $\frac{1}{T}|\langle[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}],\mathcal{W}\times_{3}^{2}\bm{X}^{*}\rangle|+\frac{A}{2},$
+where the first equation follows the fact that $(l,k,u)$-th element of $\widetilde{\bm{H}}_{D},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\widetilde{\bm{H}}_{0}]-[\bm{H}_{D}^{(p_{D})},\widetilde{\mathcal{S}}_{D},\dots,\widetilde{\mathcal{S}}_{1},\bm{H}_{0}^{(p_{0})}]$ can be decomposed into $\sum_{i=0}^{D}\bm{H}_{D}^{(p_{D})}(l,:)\widetilde{\mathcal{S}}_{D}(:,k,:)\bm{H}_{D-1}^{(p_{D-1})}\\ \cdots\bm{H}_{i+1}^{(p_{D_{i+1}})}\widetilde{\mathcal{S}}_{i+1}(:,k,:)(\widetilde{\bm{H}}_{i}-\bm{H}_{i}^{(p_{i})})\widetilde{\mathcal{S}}_{i}(:,k,:)\widetilde{\bm{H}}_{i-1}\widetilde{\mathcal{S}}_{1}(:,k,:)\widetilde{\bm{H}}_{0}(:,u)$.
 Based on the same analysis of (50), we can obtain
 $\operatorname{\mathbb{P}}\left(A\leq\frac{c_{1}\sqrt{(1+\delta_{UM})(\sum_{d=1}^{D-1}N_{d}N_{d+1})\log D}}{\sqrt{T}}\gamma\right)$ (70)
-[…]
+$\!\!\!\!\geq\!\!\!\!$ $1-2e^{-c_{2}(\sum_{d=1}^{D-1}N_{d}N_{d+1})\log D},$
 where $c_{i},i=1,2$ are constants.
 Combing (62), we arrive at
-^$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$ (66)
-[…]
+$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=1}^{D-1}N_{d}N_{d+1})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg).$
+Appendix H Error Analysis of channel estimation in the Multi-Hop IRS-Assisted MIMO Systems with special structures
+In this section, we present the error analysis of channel estimation in multi-hop IRS-assisted MIMO systems under three special structures, analogous to the single-hop case.
+$(i)$ First, when $\bm{B}_{d},d=0,\dots D$ are rank-$r_{d}$ matrices, by incorporating the covering number of low-rank matrices into Appendix G, we can derive
+$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}(N_{d}+N_{d+1})r_{d})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
+where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{D}\in\mathbb{C}^{LP\times N_{D}},\bm{B}_{0}\in\mathbb{C}^{N_{1}\times UM},\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d\in[D-1]\ \text{are unknown with }\text{rank}(\bm{B}_{d})=r_{d},d=0,\dots D,\ \text{and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$.
+Leveraging low-rank structures offers a potential avenue for improving error performance, as the recovery error may be reduced below eq.
+(14) in the main paper when the ranks $r_{d},d=0,\dots,D$ are small.
+$(ii)$ Second, when the channel matrices exhibit sparsity in the angular domain, we assume the factorized form $\bm{B}_{d}=\bm{B}_{d1}\bm{\Sigma}_{d}\bm{B}_{d2},d=0,\dots D$, where $\bm{\Sigma}_{d}\in\mathbb{C}^{F_{d}\times F_{d}}$ is a diagonal matrix with $\|\bm{\Sigma}_{d}\|_{0}=s_{d}$, and $\bm{B}_{d1}\in\mathbb{C}^{N_{d+1}\times F_{d}}$, $\bm{B}_{d2}\in\mathbb{C}^{F_{d}\times N_{d}}$.
+Following a similar approach as in Appendix E, we obtain the error bound:
+$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}=O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}s_{d}\log\big(\frac{F_{d}}{s_{d}}\big))\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
+where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{LP\times K\times UM}:\bm{B}_{d}=\bm{B}_{d1}\bm{\Sigma}_{d}\bm{B}_{d2},d=0,\dots D\ \text{where}\ \bm{\Sigma}_{d}\in\mathbb{C}^{F_{d}\times F_{d}}\ \text{is unknown}\text{with}\ \|\bm{\Sigma}_{d}\|_{0}=s_{d},\ \text{and}\ \bm{B}_{d1}\in\mathbb{C}^{N_{d+1}\times F_{d}},\bm{B}_{d2}\in\mathbb{C}^{F_{d}\times N_{d}}\ \text{are known, and}\ \mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are known}\}$.
+Note that the sparsity in the angular domain leads to a smaller recovery error compared to the non-sparse setting when $s_{d}\ll F_{d}$.
+$(iii)$ Finally, when $\mathcal{S}_{i},i\in[D]$, are unknown, introducing the covering number of $\{\mathcal{S}_{i}\}_{i=1}^{D}$ into the proof in Appendix G, we can arrive at
+$\|\widehat{\mathcal{B}}_{D}-\mathcal{B}^{\star}_{D}\|_{F}$
+$\!\!\!\!=\!\!\!\!$ $O\bigg(\frac{\gamma\sqrt{(1+\delta_{UM})(\sum_{d=0}^{D}N_{d}N_{d+1}+\sum_{d=1}^{D}KN_{d})\log D}}{(1-\delta_{UM})\sqrt{T}}\bigg),$
+where $\widehat{\mathcal{B}}_{D},\mathcal{B}^{\star}_{D}\in\{\mathcal{B}_{D}=[\bm{B}_{D},\mathcal{S}_{D},\dots,\bm{B}_{1},\mathcal{S}_{1},\bm{B}_{0}]\in\mathbb{C}^{N_{D+1}\times K\times N_{0}}:\bm{B}_{d}\in\mathbb{C}^{N_{d+1}\times N_{d}},d=0,\dots,D,\mathcal{S}_{d}\in\mathbb{C}^{N_{d}\times K\times N_{d}},d\in[D]\ \text{are unknown}\}$.
+In this analysis, we assume that all IRSs are equipped with active elements.
+However, in practice, when adjustments are required for only a subset of IRSs, denoted as $\mathcal{S}_{i}$ for $i\in\Omega$, where $\Omega\subseteq\{1,\dots,D\}$, the recovery error can be further reduced.
+Specifically, the term $\sum_{d=1}^{D}KN_{d}$ is replaced by $\sum_{i\in\Omega}KN_{i}$.